### v2

# Overcoming data challenges to measure whole-person health in electronic health records

Abstract
Objective: The allostatic load index (ALI) is a composite measure of whole-person health, which reflects the multiple interrelated physiological regulatory systems that underlie healthy functioning. Data from electronic health records (EHR) present a huge opportunity to operationalize the ALI in the learning health system, except they are prone to missingness and errors. Validation of EHR data (e.g., through chart reviews) can provide better-quality data, but realistically, only a subset of patients’ data can be validated, and most protocols do not recover missing data.
Methods: Using a representative sample of $1000$ patients from the EHR at an extensive learning health system ($100$ of whom could be validated), we propose methods to design, conduct, and analyze statistically efficient and robust studies of the ALI and healthcare utilization. With semiparametric maximum likelihood estimation methods, we robustly incorporate all available patient information (validated or unvalidated) into statistical models. Using targeted design strategies, we examine ways to select the most informative patients for validation. Incorporating clinical expertise, we devise a novel validation protocol to promote the quality and completeness of EHR data.
Results: Validation of the EHR data uncovered relatively low error rates and recovered some missing data through auxiliary information in patients’ charts. Through bespoke simulation studies based on preliminary data, residual sampling was identified as the most informative strategy for completing our validation study. Statistical models incorporating the partial validation data indicated higher odds of engaging in the healthcare system were associated with worse whole-person health (i.e., higher ALI), adjusting for age.
Conclusion: Targeted validation with an enriched protocol allowed us to ensure the quality and promote the completeness of the EHR. Findings from our validation study were incorporated into analyses as we operationalize the ALI as a whole-person health measure intended to predict healthcare utilization in the academic learning health system.
Keywords chart review  $\cdot$ data audits  $\cdot$ healthcare utilization  $\cdot$ measurement error  $\cdot$ missing data  $\cdot$ outcome-dependent sampling  $\cdot$ two-phase designs
1 Introduction
1.1 Measuring Whole-Person Health
Learning healthcare systems are transitioning from a reactive (treatment) to a proactive (prevention) philosophy. This transition could prevent future healthcare utilization through early intervention, improving health for the individual and providing cost savings to the institution. Having a standardized summary measure of health that is easy to compute and track could play a central role in disease prevention (saving lives) and reducing healthcare utilization (saving costs).
Numerous data types exist within the electronic health record (EHR), including vital signs, demographics, diagnoses, procedures, and clinical notes. Easy access to a standardized and valid summary measure of whole-person health in the EHR would be advantageous for patients, clinicians, and researchers [1, 2]. The electronic frailty index serves this purpose but is only designed for those aged 65 and older [3]. The availability of a whole-person health metric for the general adult population (age $\geq 18$) would allow for longitudinal tracking of individual and population health within a healthcare system, comparisons across geographies and hospitals, and integration of an important predictor in algorithms relevant to the enterprise (e.g., predicting readmissions).
Though individual physiological variables like systolic blood pressure measure aspects of whole-person health, they do not provide a comprehensive assessment across various organ systems. The allostatic load index (ALI) is a composite measure of whole-person health, which reflects the multiple interrelated physiological regulatory systems that underlie healthy functioning [4, 5, 6]. The ALI has been used to predict numerous health outcomes (e.g., [7, 8, 9, 10]) and measure racial disparities in health [11]. Because it is sensitive to short-term interventions, the ALI is also an appealing “surrogate” outcome to test health interventions, like lifestyle and environmental modifications.
Many ALI formulations exist. We focus on the calculation from Seeman et al. (2001) [5], which computes the ALI from ten component measurements across three body systems. As illustrated in Supplemental Figure S1, the original numeric measurements (e.g., hemoglobin A1c [HbA1c]) are discretized at clinically driven thresholds to define binary “stressor” indicators (e.g., $=1$ if hemoglobin A1c $\geq 6.5$ and $=0$ if hemoglobin A1c $<6.5$). The ALI is ultimately calculated as the sum of these ten stressors, i.e., the number of “unhealthy” measurements the patient had, with values of zero and ten representing the best and worst whole-person health, respectively.
1.2 Data Quality Challenges in Electronic Health Records Data
EHR data present a huge opportunity to operationalize the ALI in the learning health system. The volume of routinely collected data in the EHR is steadily climbing, increasing the accessibility to clinically meaningful variables. Analyzing data that come at little to no additional cost for collection is particularly appealing to biomedical researchers, leading to a surge in the utilization of EHR data in many clinical domains, including HIV and AIDS [12], genetics [13], emergency medicine [14], and therapeutic effectiveness [15].
However, despite clear advantages, EHR data are collected in the fast-paced world of clinical care, leading to missingness and errors. These quality concerns remain a significant hurdle to the large-scale adoption of EHR in healthcare research [16, 17, 18, 19]. In calculating the ALI, the original measurements may be (i) missing, making the stressor missing, or (i) non-missing but mismeasured, potentially misclassifying the stressor.
Consider a patient whose true underlying HbA1c was $6.5\%$, such that their corresponding stressor should equal one. We illustrate two scenarios for this patient’s data in Figure 1. In the first, they had HbA1c measured, but the value was erroneously transposed to $5.6\%$. Based on the EHR, this patient’s stressor for HbA1c would be misclassified as zero since $5.6\%$ falls into the healthy range. In the second, they did not have HbA1c measured, so the value is missing, and we cannot define a stressor. Simply treating missing components as zero would also induce misclassification by treating them as “healthy." We accommodate missing data by instead defining the ALI as the proportion of non-missing stressors a patient experienced.
Figure 1: Diagram of potential errors and missingness in a patient’s hemoglobin A1c (HbA1c) measurement in the electronic health record (EHR) before and after validation. In the top scenario, the patient had HbA1c measured but with error, and the correct value was recovered through validation. In the bottom scenario, the patient did not have HbA1c measured, but auxiliary information was located through validation to replace the missing component.
Ignoring errors and missingness in the EHR can hinder analyses, as biased estimates may lead to invalid inference and incorrect conclusions [20, 21]. Errors in EHR data can also reduce the accuracy of prediction models [22]. There is a particular concern for missingness and errors in derived variables, like the ALI, that combine data from many different variables [23, 24]. Analyzing the ALI computed from the EHR data “as is” may misrepresent a patient’s actual health and lead to false conclusions about the association between ALI and healthcare utilization.
For the ALI to be useful in research, it must be reliably and accurately measured. The original calculations, like Seeman et al. (2001) [5], were done in the context of clinical trials, where all patients must have all measurements and data validation protocols are likely in place to ensure quality. Similarly, validation, wherein the extracted EHR data are compared to original source documents (like the electronic chart) provides insights into data quality. Complete data validation (i.e., for all patients and variables) would ensure the integrity of EHR data but is overly expensive and unattainable in practice.
Fortunately, partially validating EHR data can help identify issues and pave the way for corrections. It provides a cost-effective alternative to complete data validation and is a promising approach to gauge the quality of observational data sources [25, 26, 27]. These partial validation studies are a special case of two-phase designs [28, 29] (Figure 2A). Phase I includes variables available for all $N$ patients, which can be used to select a subsample of $n$ patients ($n<N$) for validation in Phase II.
Figure 2: Diagram of two-phase designs with $N$ patients extracted from the electronic health records (EHR) in phase I and $n$ patients chosen for validation via chart review in phase II. The least informative strategy to select the validation study is A) simple random sampling. More informative targeted strategies include B) case-control sampling, C) balanced case-control sampling using discretized ALI, D) optimal stratified sampling using discretized ALI, and E) residual sampling based on the naive model. Within strata, the darker shaded areas indicate which patients would be selected.
Validated data are usually expected to be more correct (i.e., less error-prone) but not more complete (i.e., less missingness). To calculate the ALI from EHR data, we needed validation to do both. Herein, we devise a novel validation protocol that incorporates a “roadmap” created by clinical experts to recover missing data created by clinical experts using auxiliary health information (e.g., related diagnoses).
System Component Threshold Auxiliary Information
Cardiovascular Systolic Blood Pressure $>140$ Hypertension
Diastolic Blood Pressure $>90$ Hypertension
Metabolic Body Mass Index $>30$ Obesity; Morbid Obesity;
Grade I, II or III Obesity
Triglycerides $\geq 150$ Hypertriglyceridemia
Total Cholesterol $\geq 200$ Hypercholesterolemia
Inflammation C-Reactive Protein $\geq 10$ Sepsis; Infection; Auto-Immune
Inflammatory Syndrome
Hemoglobin A1C $\geq 6.5$ Diabetes; Impaired
Glycemic Control
Serum Albumin $\geq 3.5$ (None Given)
Creatinine Clearance $<110$ (Males) Renal Failure;
$<100$ (Females) Insufficiency; Acute
Kidney Injury;
Chronic Renal Failure
Homocysteine $>50$ Hyperhomocysteinemia;
Vitamin deficiency
Table 1: Ten component stressors of the allostatic load index were defined by discretizing measurements across three body systems at clinically driven thresholds. If the component was missing from the extracted EHR data, auditors searched for auxiliary information in the patient’s medical chart in Epic. If the auxiliary information was present, the component was treated as a “yes” and included in the patient’s validated ALI.
1.3 Overview
While very insightful, partial validation studies are expensive, time-intensive undertakings. Using a representative sample of $1000$ patients from the EHR at an extensive academic learning health system ($100$ of whom could be validated), we propose novel methods to design, conduct, and analyze statistically efficient and robust studies of whole-person health (ALI) and healthcare utilization. Using advanced statistics and bioinformatics methods, we seize three key opportunities to make the most of our investment: (i) robust and efficient statistical modeling, (ii) enriched validation protocol, and (iii) targeted validation study design.
Statement of Significance
Problem Standardized and valid summary measures of current health, i.e., whole-person health, are needed in the EHR to make patient- and institution-level decisions. However, routinely collected data are prone to missingness and errors.
What Is Already Known Validation, wherein the extracted EHR data are compared to the original clinical source documents (like the electronic chart), provides insights into data quality, but validating all patients is not feasible, and existing protocols can only correct for errors, not missingness.
What This Paper Adds Building on the statistics and bioinformatics literature, we 1) devise a robust and efficient model for partially validated, error-prone EHR data; 2) propose a targeted validation study design to identify the most informative patients for our model; and 3) develop an enriched validation protocol that incorporates clinical expertise to improve the quality and completeness of EHR data.
2 Methods
2.1 Model and Data
This study was approved by the Institutional Review Board at Wake Forest University School of Medicine. We are interested in modeling healthcare utilization ($Y\in\{0,1\}$), defined as a binary indicator of having at least one hospitalization or emergency department (ED) visit in the $2$-year study period. The exposure of interest is the ALI ($X\in[0,1]$), defined as the proportion of stressors a patient experienced. Further, we control for patient age at first encounter ($Z\in\mathcal{R}$).
These variables are assumed to be related through the usual logistic regression model for the log-odds that $Y=1$ (engaged in the healthcare system) given $X$ and $Z$. A sample of $N$ patients was extracted from the EHR to estimate these log odds ratios (ORs), denoted ${\boldsymbol{\beta}}$. Due to missingness and errors in the EHR, only an error-prone surrogate measure for $X$, denoted by $X^{*}$ ($X^{*}\in[0,1]$), is available. We compute $X^{*}$ as the proportion of non-missing stressors the patient experienced, according to the EHR data. Age $Z$ is fully observed and expected to be error-free.
To obtain consistent estimates for the desired healthcare utilization model, a two-phase design is planned (Figure 2A). In Phase I, the outcome $Y$ and surrogate exposure $X^{*}$ will be collected from the EHR for all $N$ patients. In Phase II, these data will be augmented with the validated exposure $X$ from a chart review for a subset of $n$ patients ($n<N$). To distinguish validated and unvalidated patients, let $V_{i}$ be a validation indicator, where $V_{i}=1$ if the $i$th patient is validated, $i\in\{1,\dots,N\}$, and $V_{i}=0$ otherwise.
Incorporating data from Phases I and II, along with the validation indicator, the joint distribution of a complete observation is assumed to be
$\Pr(V,X^{*},Y,X,Z)$ $=\Pr(V\mid Y,X^{*},Z){\Pr}_{{\boldsymbol{\beta}}}(Y\mid X,Z)\Pr(X\mid X^{*},Z)\Pr(X^{*},Z),$ (1)
where $\Pr(V\mid Y,X^{*},Z)$ is the validation sampling probability; ${\Pr}_{{\boldsymbol{\beta}}}(Y\mid X,Z)$ corresponds to the logistic regression model; $\Pr(X\mid X^{*},Z)$ is the exposure error mechanism; and $\Pr(X^{*},Z)$ is the joint probability density or mass function of $X^{*}$ and $Z$. Two assumptions are made in defining (1). First, selection for validation (i.e., $V$) is assumed to depend only on Phase I variables ($Y$, $X^{*}$, $Z$), so the validated exposure $X$ is missing at random (MAR) for unvalidated patients [30]. Second, the exposure $X^{*}$ from the EHR is assumed to be a surrogate for the validated exposure $X$, such that $Y$ and $X^{*}$ are conditionally independent given $X$. In other words, once $X$ is known, $X^{*}$ is assumed not to contribute anything further about the relationship with $Y$. While other ways of factoring (1) are possible, this factorization avoids making an additional assumption about $\Pr(X^{*},Z)$.
2.2 Robust Statistical Model for Partially Validated Data
The observed-data log-likelihood function (hereafter, log-likelihood function) for the log ORs ${\boldsymbol{\beta}}$ is defined based on the joint distribution in (1), but first, two simplifications can be made. First, the validation sampling probability $\Pr(V|Y,X^{*},Z)$ can be omitted from the log-likelihood for ${\boldsymbol{\beta}}$ because the Phase II variables are MAR. Second, the joint probability $\Pr(X^{*},Z)$ can be ignored because it will drop out as the log-likelihood function is maximized with respect to ${\boldsymbol{\beta}}$ since $X^{*}$ and $Z$ are fully observed.
Observations for all $N$ patients in the study are assumed to be independent and identically distributed (i.i.d.) following (1). Now, the log-likelihood function for ${\boldsymbol{\beta}}$ is proportional to the following:
$\sum_{i=1}^{N}V_{i}\left[\log\left\{{\Pr}_{{\boldsymbol{\beta}}}(Y_{i}\mid X_{i},Z_{i})\right\}+\log\left\{\Pr(X_{i}\mid X_{i}^{*},Z_{i})\right\}\right]$
$+\sum_{i=1}^{N}(1-V_{i})\log\left\{\sum_{x=0}^{1}{\Pr}_{{\boldsymbol{\beta}}}(Y_{i}\mid x,Z_{i})\Pr(x\mid X_{i}^{*},Z_{i})\right\}.$ (2)
Notice that unvalidated patients contribute their incomplete information to this log-likelihood by summing over the missing validated exposure $X$:
$\sum_{x=0}^{1}{\Pr}_{{\boldsymbol{\beta}}}(Y_{i}\mid x,Z_{i})\Pr(x\mid X_{i}^{*},Z_{i})$ $\propto\sum_{x=0}^{1}{\Pr}_{{\boldsymbol{\beta}}}(Y_{i},x,Z_{i},X_{i}^{*})={\Pr}_{{\boldsymbol{\beta}}}(Y_{i},Z_{i},X_{i}^{*}),$ (3)
remembering that $X$ takes on values $x\in\{0,0.1,\dots,0.9,1\}$ here.
Still, a model for the exposure error mechanism $\Pr(X\mid X^{*},Z)$ is needed to estimate the analysis model parameters since $X$ is only observed for the subset of validated patients. Since the exposure $X$, representing the patient’s ALI, is bounded between zero and one (inclusive), there are not many flexible, parametric choices to model the distribution of $X$ given ($X^{*},Z$). Thus, a nonparametric estimator is preferable, allowing the validated exposure $X$ to take on any values and arbitrarily depend on $X^{*}$ and $Z$.
Lotspeich et al. (2022) [31] proposed a semiparametric approach to logistic regression with outcome and exposure errors, which nonparametrically (i.e., without assuming a particular distribution or structure) estimates the exposure error mechanism as described. The current setting is a special case with only exposure error, and a brief outline is provided in the Supplemental Materials. For full technical details on the semiparametric sieve maximum likelihood estimators (SMLEs), see Lotspeich et al. (2022) [31].
2.3 Targeted Designs to Select Informative Patients for Validation
As a baseline, consider the simple random sampling (SRS) design, wherein $n$ of the $N$ patients are chosen for validation with equal probability. This strategy ignores all available information, making it the most uninformative. For full-likelihood approaches like the SMLE, any information available for all patients can – and should – be used to design the validation study. Designing based on fully observed information ensures that the MAR assumption holds for the downstream analysis while allowing for many targeted strategies.
When designing a validation study to promote statistical efficiency of the parameter estimates (i.e., to minimize the variance in the estimated log OR on ALI), we start at the end. We begin with the planned logistic regression model of $Y$ given $X$ and $Z$, and work backward to determine which patients will be most informative to it. For this reason, the “best” designs vary across different estimators, with model-based estimators, like the SMLEs, offering the most design flexibility. Other estimators that rely on the design–rather than a model–to handle the missingness, like inverse probability weighting, require nonzero sampling probability for all patients, but model-based estimators do not.
We considered stratified and rank-based validation study designs. Multiple validation waves were conducted to allow the protocol and study design to adapt as information accumulated, which is becoming common in this space [32, 26, 33]. Ultimately, simulations based on preliminary validation data (see Section 3.3.3) guided our selection of the targeted design used in the application.
Given our binary outcome, the case-control (CC) design is a natural option here, wherein $n/2$ patients each are chosen from subgroups with $Y=0$ and $Y=1$ (Figure 2B). The balanced case-control (BCC) design further stratifies on covariate information, subsequently choosing equal (i.e., balanced) numbers from each stratum (Figure 2C). When the covariates are continuous, they must be discretized to form strata. For instance, we could use $X_{D}^{*}$, a discretized version of $X^{*}$, to define strata of patients with error-prone ALI above and below the average. Then, $n/4$ patients are validated from each of the four $(Y,X_{D}^{*})$ strata. Importantly, we can still fit the model using the original continuous values.
While the CC and BCC designs are intuitive and can offer some efficiency gains over SRS, there are slightly more complicated designs that are guaranteed to be theoretically “optimal,” meaning that they choose the number of patients to validate from each stratum to achieve the lowest possible variance for the log OR on ALI. Many have been proposed to select patients for the collection of “expensive” binary exposures, like biomarker measurement or validation via chart review, for a variety of analytical approaches to modeling binary outcomes [34, 35, 36, 37, 29, 38]. Recently, Lotspeich et al. (2024) [33] proposed a design for optimal validation of secondary use data for maximum likelihood estimation with binary outcome and exposure misclassification (Figure 2D), which could be applied here. However, we would achieve the lowest possible variance for the discretized version $X_{D}^{*}$ rather than $X^{*}$. Increasing the number of unique values of $X_{D}^{*}$ (i.e., defining more strata) can help, but information will still be lost.
Still, designs that rank patients rather than stratifying them offer improved granularity and sizable efficiency gains. For example, extreme tail sampling (ETS) designs focus on the extremes (i.e., tails) of the distribution, ranking the data by some value of interest and targeting patients with the smallest or largest values. To identify patients who are most informative not just about ALI ($X^{*}$ and $X$) but about its relationship to healthcare utilization ($Y$), residual sampling (RS) is preferred [39]. With measurement error, we begin by fitting the “naive” analysis model for $Y$ (i.e., using $X^{*}$ instead of $X$) to obtain estimates for the log ORs $\hat{\boldsymbol{\beta}}^{*}$. Then, residuals for the naive model are computed for each patient $i$ ($i=1,\dots,N$) as
$r_{i}^{*}$ $=Y_{i}-{\Pr}_{\hat{\boldsymbol{\beta}}^{*}}(Y_{i}=1|X^{*}_{i},Z_{i})$
$=Y_{i}-\frac{1}{1+\exp\left\{-\left(\beta_{0}^{*}+\beta_{1}^{*}X_{i}^{*}+{\boldsymbol{\beta}}_{2}^{*}Z_{i}\right)\right\}},$ (4)
and used to rank the $N$ patients in the study. With logistic regression, the largest residuals $r_{i}^{*}$ (in magnitude) correspond to observations where the observed value $Y_{i}$ had a low predicted probability (Figure 2E). The auditDesignR package, available on GitHub at www.github.com/sarahlotspeich/auditDesignR, contains R functions to sample from all of these designs.
2.4 Enriched Validation Protocol to Reduce Errors and Missingness
The chart reviews to validate the EHR data were conducted over $6$ months (February to July 2024) by a team of four clinical research assistants and technicians in the Wake Forest University School of Medicine (hereafter, “auditors”). These skilled auditors were chosen for their familiarity with Epic, the institution’s electronic charting software, and their demonstrated experience in the clinical research environment.
2.4.1 Identifying Sources of Auxiliary Information
Before the chart reviews, our clinical collaborators developed the “audit roadmap” (Table 1). While it is unlikely that we would identify the original missing measurements, we harnessed the clinical expertise of our team to identify sources of auxiliary information about the missing discretized components. The roadmap suggested where auditors could look in Epic to locate supplemental information about missing values.
For example, if a patient were missing HbA1c, the auditors would search their chart for a diagnosis of diabetes or impaired glycemic control. If either were present, the patient’s diagnosis would suggest their missing HbA1c value would have exceeded the clinically meaningful threshold of $6.5\%$. Thus, their validated ALI $X$ would include one point for HbA1C. If both were absent, the patient’s validated ALI would continue to exclude this still-missing component.
Many components, like C-reactive protein and HbA1c, had multiple sources of auxiliary information, while others, like triglycerides and total cholesterol, had just one. Only serum albumin, a blood test to measure liver and kidney function, did not have a supplemental diagnosis.
2.4.2 Validated Data Collection
The EHR data were extracted by the Wake Forest Clinical and Translational Science Institute and securely received by the analytical team. In preparation for validation, the analytical team transformed the data from the original wide format (with one row per patient encounter and columns for different variables) to a long format (with one row per patient encounter per variable). This transformation streamlined the data structure for the auditors, with only one value to validate per row (Figure 3).
Figure 3: Illustrative example of how the original extracted EHR data for patients’ vitals were transformed from a wide format (with one row per patient encounter and columns for different variables) to a long format (with one row per patient encounter per variable), merged with the audit roadmap, and augmented with empty columns for auditors to enter their findings. The same was done with patients’ lab data.
Before sending the data to the auditors, the roadmap was incorporated, and an empty column was added for the “reviewed value” (i.e., the one found in Epic). Depending on the scenario, the auditors would enter different information for the reviewed value, and each data point was assigned an audit finding (flow chart in Figure 4). Three findings were possible for non-missing components in the EHR.
  1. 1.
    Extracted value correct: If the component matched Epic, the auditor would enter the original value again.
  2. 2.
    Extracted value incorrect: If the component did not match Epic, the auditor would enter the new value found in Epic.
  3. 3.
    Extracted value not found: If the component could not be found in Epic, the auditor would enter “Not Found.”
Two findings were possible for missing components.
  1. 4.
    Auxiliary information found: If auxiliary information from the audit roadmap were found in Epic, the auditor would enter the auxiliary information.
  2. 5.
    Auxiliary information not found: If auxiliary information from the audit roadmap were not found in Epic, the auditor would enter “Not Found.”
Figure 4: Flow chart of instructions given to auditors for the chart reviews (Steps 1–3) and the corresponding audit finding assigned by the analysts (Step 4). Column names refer to those in Figure 3.
3 Results
3.1 Cohort Description
We obtained a representative sample of $N=1000$ patients from the EHR at Atrium Health Wake Forest Baptist Hospital in Winston-Salem, North Carolina. To be considered, patients needed to (i) be $18-65$ years old and (ii) initiate their first primary care outpatient encounter in the health system between March 11, 2018, and March 10, 2020 (to avoid data challenges and biases associated with the COVID-19 pandemic). Outpatient encounters were chosen to represent the typical patient voluntarily engaging in care.
Our data contained demographics, vitals, lab values, discharge diagnoses, and problem list values dating $3$ years before the first encounter. Key variables included the measurements for the ten components of the ALI (Table 1), age at first encounter (hereafter, “age”), and an indicator of having at least one hospitalization or ED visit during the $2$-year study period. Since most patients had multiple encounters, the ALI components were based on mean measurements.
Sampled patients had a median age of $48$ years and a median income of \$52,320. They were $61\%$ female, $72\%$ White, $6\%$ Hispanic, and $73\%$ homeowners. There were $318$ patients ($32\%$) with at least one hospitalization or ED visit. Based on the original unvalidated EHR data, the distribution of error-prone ALI was right-skewed (Supplemental Figure S2). With a median of $0.33$ (interquartile range $=0.17,0.50$), on average, patients in our study experienced $1$ out of every $3$ non-missing stressors on their whole-person health.
3.2 Baseline Data Quality and Naive Model Estimates
As expected, the extracted EHR data had a large amount of missingness. In particular, rare labs for inflammation components were missing for $>95\%$ of patients. Labs for the metabolic components were missing for $21\%$ of patients, and measurements for the cardiovascular components were non-missing for basically everyone (Supplemental Figure S3A). The most common patterns of missing data are displayed in Supplemental Figure S4. Fortunately, most patients ($78\%$) had at least $5$ non-missing ALI components (Supplemental Figure S3B).
Using these ALI values to fit the “naive” healthcare utilization model, we estimate the following. First, for every $0.1$-point (or one component) increase in ALI, the odds of engaging in the healthcare system are expected to increase by $10\%$ (OR $=1.10$, 95% confidence interval [95% CI]: $1.03$, $1.18$), adjusting for age. Second, the expected odds of engaging in the healthcare system for an $18$-year-old patient with an ALI of 00 are $0.25$ (95% CI: $0.18$, $0.35$). Third, for every $10$ years older a patient was at their first encounter, their odds of engaging in the healthcare system are expected to increase by $11\%$ (OR $=1.11$, 95% CI: $1.00,1.23$), adjusting for ALI. Thus, the naive model suggests that worse whole-person health leads to higher expected odds of hospitalization or emergency department visits.
3.3 Three Waves of Data Validation Via Chart Review
While the naive model aligns with the existing literature, we wanted to ensure that the missingness and potential errors in the EHR data did not lead to spurious conclusions. Thus, we secured funding to validate data for $n=100$ of the $N=1000$ patients, whose selection and review were broken down into three waves: (i) Pilot, (ii) Wave I, and (iii) Wave II.
3.3.1 Pilot Wave: Testing the Audit Protocol and Establishing Inter-Auditor Agreement
The first $n_{0}=4$ patients were selected for chart review via the BCC design introduced in Section 2.3. Specifically, patients were split into four strata based on their outcome and error-prone ALI, discretized at the median:
  • (i)
    Engaged in the healthcare system and low error-prone ALI ($Y=1$, $X^{*}\leq 0.33$),
  • (ii)
    Engaged in the healthcare system and high error-prone ALI ($Y=1$, $X^{*}>0.33$),
  • (iii)
    Did not engage in the healthcare system and low error-prone ALI ($Y=0$, $X^{*}\leq 0.33$), and
  • (iv)
    Did not engage in the healthcare system and high error-prone ALI ($Y=0$, $X^{*}>0.33$).
Then, $1$ patient was randomly chosen for validation from each stratum.
The auditors followed the protocol in Section 2.4 to (i) verify the non-missing values in the EHR data or note any discrepancies and (ii) locate auxiliary information for missing values. Following the completion of the Pilot wave, the auditors and analysts met to discuss any modifications that would be helpful before proceeding. For example, the third audit finding of “extracted value not found” was not one of the original options, and the team discussed how these data points should be recorded. Adding an optional “Notes” column, where auditors could include any additional information they uncovered, also resulted from this discussion. The Pilot and Wave I findings were combined to interpret data quality (see Section 3.3.2).
We started with two auditors on our team. Each reviewed the four patients chosen for the Pilot, allowing us to calculate inter-auditor agreement in this doubly-validated sample. Later, two new auditors joined the team, each reviewing these four patients’ records as well. Ultimately, agreement in the four auditors’ findings was calculated using Fleiss’ Kappa [40], as detailed in the Supplemental Materials. We found high agreement between the auditors’ findings for lab ($\kappa=0.90$) and vitals ($\kappa=0.81$) variables.
3.3.2 Wave I: Collecting Preliminary Audit Data
The next $n_{1}=48$ patients validated were also selected via the BCC design from Section 3.3.1. Since we did not have any prior information about the quality of the EHR data, we sampled more conservatively here to collect preliminary information for a more extreme, targeted design in Wave II. Multi-wave validation schemes like this one are appealing because they balance learning about our data with strategically designing for them [32, 26]. Even if prior information is available (e.g., from a previous audit), validating in multiple waves allows us to relearn how our data may have changed before using all our resources to validate them [33].
[…]
Figure 5: Auditors’ findings from the validation of A) $52$ patients’ data in the Pilot + Wave I, B) $48$ patients’ data in Wave II, and C) $100$ patients’ data in all waves combined. These findings refer to the original numeric measurements (before discretizing them into allostatic load index components), and there could be multiple per patient per variable.
Fewer vitals than labs were found to match Epic ($59\%$ versus $81\%$), with $31\%$ not found. Through validation, we discovered that the EHR data extraction algorithm expected every patient encounter to have vitals. For encounters where vitals were not taken (e.g., physical therapy or optometry visits), values were imputed from other encounters; these imputed values could not be found in Epic by the auditors. Thus, we uncovered an area for potential improvement in the current EHR data extraction protocol.
Figure 6: Comparison of the discretized allostatic load index components from the unvalidated EHR data to A) $52$ patients’ data in the Pilot + Wave I, B) $48$ patients’ data in Wave II, and C) $100$ patients’ data in all waves combined.
Incorporating the audit roadmap into the validation protocol improved the completeness of the data, as the number of missing values decreased for nearly all components of the ALI (Supplemental Figure S3A). Homocysteine saw the biggest improvement, with auxiliary information located for $23$ of the $52$ validated patients ($44\%$). The number of non-missing ALI components per patient increased from $6.6$ to $7.5$, on average. Overall, $48$ of $177$ missing components ($27\%$) were recovered through validation (Figure 6A).
Incorporating the Pilot and Wave I, we re-fit the healthcare utilization model using the SMLE with the first $52$ validated and $948$ unvalidated patients’ data (Table 2). The association between whole-person health and healthcare utilization appeared even stronger, with an estimated 21% increase in odds for each $0.1$-point increase in ALI (OR $=1.21$, 95% CI: $1.17,1.25$), adjusting for age. For an 18-year-old patient with an ALI of 00, the expected odds of engaging in the healthcare system decreased to $0.17$ (95% CI: $0.12$, $0.26$). The adjusted OR for age was relatively unchanged (OR $=1.10$, 95% CI: $1.01,1.21$). Thus, the preliminary SMLE model reaffirms that worse whole-person health is associated with higher expected odds of engaging in the healthcare system.
Odds/Odds
Coefficient Ratio Estimate 95% CI
Naive Analysis (Unvalidated Data)
Intercept $0.251$ $(0.176,0.357)$
ALI $1.099$ $(1.028,1.175)$
Age $1.108$ $(1.000,1.228)$
Preliminary SMLE (Pilot + Wave I Validated Data)
Intercept $0.173$ $(0.115,0.262)$
ALI $1.207$ $(1.167,1.247)$
Age $1.104$ $(1.007,1.211)$
Final SMLE (All Waves Validated Data)
Intercept $0.225$ $(0.156,0.327)$
ALI $1.120$ $(1.088,1.154)$
Age $1.105$ $(1.018,1.200)$
Table 2: Estimates and 95% confidence intervals for coefficients from the naive analysis (logistic regression, no validated data), preliminary sieve maximum likelihood estimator (SMLE, validated data on $52$ patients), and final SMLE (validated data on $100$ patients) models.
3.3.3 Simulations: Using Preliminary Audit Findings to Design Wave II
Before validating the remaining patients, we designed bespoke simulation studies based on the preliminary audit data. These simulations allowed us to explore how potential designs compared under different scenarios for the true underlying data quality. In particular, the design offering the best statistical efficiency for the OR on ALI was desired. Focusing here on the takeaways and our decision-making, details like the data-generating mechanisms can be found in the Supplemental Materials. The five designs described in Section 2.3 were considered: SRS, CC, BCC, OPT, and RS.
Under varied error rates, the targeted validation study designs offered empirically unbiased estimates and higher efficiency (i.e., smaller variance) than the uninformative SRS strategy. The RS design provided the best efficiency under extra low to moderate error settings; the preliminary validation data indicated that we had a perfect TPR $=1.00$ and negligible FPR $<0.01$.
The SMLE $\hat{\beta}_{1}$ was empirically unbiased under full/high data recovery rates ($90-100\%$ missing data recovered through validation) and different validation study designs. While biased for moderate recovery or lower ($\leq 50\%$), the SMLE was still closer to the truth than the naive estimate was (i.e., based on the unvalidated EHR data). The RS design offered the highest efficiency, regardless of how much missing data was recovered. Thus, based on extensive simulations considering different error rates and data recovery scenarios, we concluded that the RS design was the most informative way to select the remaining $n_{2}=48$ patients for validation in Wave II.
3.3.4 Wave II: Implementing the Targeted Validation Study Design
For the RS design, each unvalidated patient’s residual was calculated based on the naive model estimates, following (4), as
$r_{i}^{*}=Y_{i}-\frac{1}{1+\exp\left\{-\left(-1.383+0.945X_{i}^{*}+0.103Z_{i}\right)\right\}}.$
Then, patients were ranked by $r_{i}^{*}$, and the $24$ patients with the smallest residuals and the $24$ patients with the largest residuals were selected for validation. Here, the RS design prioritized (i) patients with zero hospitalizations or ED visits but high predicted probabilities of having at least one and (ii) patients with at least one hospitalization or ED visit but high predicted probabilities of having zero. That is, patients whose observed $Y_{i}$ aligned well with the naive model’s prediction were not considered to be very informative and thus were deprioritized for validation.
A total of $3867$ data points were validated in Wave II ($718$ lab and $3149$ vitals variables), which was fewer than in the Pilot and Wave I. Given that similar numbers of patients were validated, this difference was explained by fewer patient encounters, with between $16$ and $453$ validated data points per patient (median $=65.5$). Overall, the estimated error rates for non-missing components (TPR $=99\%$, FPR $=3\%$) and recovery rate for missing data ($17\%$) in Wave II closely resembled the preliminary ones (Figure 6B). Still, some auditors’ findings differed (Figure 5B). For example, more vitals were found to match Epic ($65\%$). Similar amounts of missing data were recovered using the audit roadmap, spread across all lab variables. Additional findings are in Supplemental Table S1.
3.4 Post-Validation Model of Healthcare Utilization After Validation
Combined findings from all waves indicated that the non-missing EHR data were of excellent quality (TPR $=99.5\%$, FPR $=2\%$), and auxiliary information could be located for $21\%$ of missing components (Figures 5C and 6C). Detailed findings are in Supplemental Table S2. We pooled validated data for the $n=100$ total validated patients with the unvalidated on all $N=1000$ patients from the original EHR data to estimate the SMLE once more (Table 2).
All final model estimates fell between the naive analysis and preliminary SMLEs. For each $0.1$-point increase in ALI, a $12\%$ increase in odds of engaging in the healthcare system was expected (OR $=1.12$, 95% CI: $1.09,1.15$), adjusting for age. The expected odds of engaging in the healthcare system for an 18-year-old patient with an ALI of 00 was $0.23$ (95% CI: $0.16$, $0.33$). Again, older patients were expected to be more likely to engage in the healthcare system, controlling for ALI (OR $=1.11$, 95% CI: $1.02,1.20$). Additional technical details, including R code, are available in the Supplemental Materials.
Ultimately, all three models agreed that worse whole-person health (i.e., higher ALI) was associated with a higher likelihood of being hospitalized or visiting the emergency department, adjusting for age. Our chart reviews uncovered very high quality of the non-missing EHR data and located auxiliary information for some missing data. Without validation, knowing whether we can trust the EHR data would not have been possible. Thus, while the validated and unvalidated models led to the same conclusions here, validating EHR data before using them in research is critical.
4 Discussion
Predicting healthcare utilization is a priority in the learning health system, with important implications at both the patient- and institution-levels. Establishing standardized, valid summary measures of whole-person health, like the ALI, associated with utilization and can scale up to the entire institution is an integral step. While EHR data contain variables to calculate these measures, missing data is a key challenge, and non-missing data are still prone to errors.
Data validation or auditing, often through chart reviews, can gauge data quality, but this time-intensive process is usually only possible for a small subset of patients’ data. Also, existing validation protocols do not overcome missingness. Herein, we propose novel methods to maximize information gained from a partial validation study of $100$ patients from an EHR dataset of $1000$ through the statistical model, validation protocol, and targeted strategy to select patients.
The proposed analytical and design strategies were devised to ensure robust and efficient estimates. Specifically, the statistical model avoids making restrictive assumptions about the error mechanism, relating the unvalidated and validated ALI, while incorporating all available information to achieve precise estimates (i.e., low variance). Moreover, the targeted design strategies are tailored to the preliminary validation data and chosen statistical model, selecting patients most informative for the clinical question. Our enriched validation protocol builds on those in the literature (e.g., [41]) by incorporating a clinically-driven “roadmap” to locate auxiliary information about missing values.
While our roadmap helped recover some missing ALI components, the clinical guidelines were only defined for missing values that were presumed to be a “yes.” If auditors were given additional guidance on what auxiliary information might mean that a missing value is a “no,” the missingness could be further reduced. Others have described missingness in the EHR as selection bias and sought to correct for it using approaches like inverse probability weighting [42]. Applying such approaches to our post-validation data may reduce the impact of residual missingness.
Further, EHR data are susceptible to “informative presence bias,” wherein the process through which patients engage in the healthcare system can bias associations [43, 44, 45]. For example, patients seeking care more often may have more diagnoses. Our decision to adapt the Seeman et al. (2001) [5] definition of the ALI to be the proportion of non-missing components could be considered a summary measure. In a recent review, Sisk et al. (2021) [46] found summary measures to be the most common approach to handling informative presence bias. We also used the enriched chart reviews to better understand the missingness.
There are several interesting directions for future work. First, patients could have multiple labs and vitals during the study period. The averages of these values were used to calculate the ALI, and a longitudinal approach to modeling healthcare utilization wherein patients’ ALI can vary over time might provide more information. Second, auditors recorded their findings in a Microsoft Excel workbook. An electronic REDCap [47] interface could be built following published guidelines [48] for ease of use and accuracy. Third, one could use other machine learning techniques (e.g., random forest classification) and explore targeted validation strategies to improve predictive ability (e.g., accuracy) rather than statistical efficiency [49, 50].
5 Conclusion
EHR data provide an exciting opportunity to implement key measures, like whole-person health scores, in the learning health system. However, errors and missingness in these routinely collected data pose challenges in verifying and operationalizing such measures. Validation of even a subset of patients’ data is time- and cost-intensive but provides crucial insights into data quality. Statistical and biomedical informatics methods can maximize information gained from such partial validation.
Acknowledgments
The authors gratefully acknowledge Wake Forest University and Wake Forest University School of Medicine for an intercampus collaborative grant that supported this work. Computations were performed using the Wake Forest University (WFU) High Performance Computing Facility, a centrally managed computational resource available to WFU researchers including faculty, staff, students, and collaborators.
Supplementary Materials

### v4

@@
-Overcoming data challenges to measure whole-person health in electronic health records
+Overcoming data challenges through enriched validation and targeted sampling to measure whole-person health in electronic health records
 Abstract
-Objective: The allostatic load index (ALI) is a composite measure of whole-person health, which reflects the multiple interrelated physiological regulatory systems that underlie healthy functioning.
-Data from electronic health records (EHR) present a huge opportunity to operationalize the ALI in the learning health system, except they are prone to missingness and errors.
-Validation of EHR data (e.g., through chart reviews) can provide better-quality data, but realistically, only a subset of patients’ data can be validated, and most protocols do not recover missing data.
-Methods: Using a representative sample of $1000$ patients from the EHR at an extensive learning health system ($100$ of whom could be validated), we propose methods to design, conduct, and analyze statistically efficient and robust studies of the ALI and healthcare utilization.
-With semiparametric maximum likelihood estimation methods, we robustly incorporate all available patient information (validated or unvalidated) into statistical models.
+Objective: The allostatic load index (ALI) is a 10-component composite measure of whole-person health, which reflects the multiple interrelated physiological regulatory systems that underlie healthy functioning.
+Data from electronic health records (EHR) present a huge opportunity to operationalize the ALI in learning health systems; however, these data are prone to missingness and errors.
+Validation (e.g., through chart reviews) can provide better-quality data, but realistically, only a subset of patients’ data can be validated, and most protocols do not recover missing data.
+Methods: Using a representative sample of $1000$ patients from the EHR at an extensive learning health system ($100$ of whom could be validated), we propose methods to design, conduct, and analyze statistically efficient and robust studies of ALI and healthcare utilization.
+Employing semiparametric maximum likelihood estimation, we robustly incorporate all available patient information into statistical models.
 Using targeted design strategies, we examine ways to select the most informative patients for validation.
-Incorporating clinical expertise, we devise a novel validation protocol to promote the quality and completeness of EHR data.
-Results: Validation of the EHR data uncovered relatively low error rates and recovered some missing data through auxiliary information in patients’ charts.
-Through bespoke simulation studies based on preliminary data, residual sampling was identified as the most informative strategy for completing our validation study.
-Statistical models incorporating the partial validation data indicated higher odds of engaging in the healthcare system were associated with worse whole-person health (i.e., higher ALI), adjusting for age.
-Conclusion: Targeted validation with an enriched protocol allowed us to ensure the quality and promote the completeness of the EHR.
-Findings from our validation study were incorporated into analyses as we operationalize the ALI as a whole-person health measure intended to predict healthcare utilization in the academic learning health system.
-Keywords chart review  $\cdot$ data audits  $\cdot$ healthcare utilization  $\cdot$ measurement error  $\cdot$ missing data  $\cdot$ outcome-dependent sampling  $\cdot$ two-phase designs
+Incorporating clinical expertise, we devise a novel validation protocol to promote EHR data quality and completeness.
+Results: Chart reviews uncovered few errors ($99\%$ matched source documents) and recovered some missing data through auxiliary information in patients’ charts.
+On average, validation increased the number of non-missing ALI components per patient from $6$ to $7$.
+Through simulations based on preliminary data, residual sampling was identified as the most informative strategy for completing our validation study.
+Incorporating validation data, statistical models indicated that worse whole-person health (higher ALI) was associated with higher odds of engaging in the healthcare system, adjusting for age.
+Conclusion: Targeted validation with an enriched protocol enabled us to ensure the quality and promote the completeness of the EHR.
+Findings from our validation study were incorporated into analyses as we operationalize the ALI as a scalable whole-person health measure predictive of healthcare utilization in the learning health system.
+Keywords chart review  $\cdot$ computable phenotype  $\cdot$ data audits  $\cdot$ healthcare utilization  $\cdot$ measurement error  $\cdot$ missing data  $\cdot$ residual sampling  $\cdot$ two-phase designs
 1 Introduction
 1.1 Measuring Whole-Person Health
-Learning healthcare systems are transitioning from a reactive (treatment) to a proactive (prevention) philosophy.
+The advent of learning health systems [1, 2] signals a transition from a reactive (treatment-based) to a proactive (prevention-based) philosophy in healthcare.
 This transition could prevent future healthcare utilization through early intervention, improving health for the individual and providing cost savings to the institution.
-Having a standardized summary measure of health that is easy to compute and track could play a central role in disease prevention (saving lives) and reducing healthcare utilization (saving costs).
+Having a standardized summary measure of health that is simple to compute and track could play a central role in disease prevention (saving lives) and reducing healthcare utilization (saving costs).
 Numerous data types exist within the electronic health record (EHR), including vital signs, demographics, diagnoses, procedures, and clinical notes.
 Easy access to a standardized and valid summary measure of whole-person health in the EHR would be advantageous for patients, clinicians, and researchers [3, 4].
-The electronic frailty index serves this purpose but is only designed for those aged 65 and older [3].
-The availability of a whole-person health metric for the general adult population (age $\geq 18$) would allow for longitudinal tracking of individual and population health within a healthcare system, comparisons across geographies and hospitals, and integration of an important predictor in algorithms relevant to the enterprise (e.g., predicting readmissions).
-Though individual physiological variables like systolic blood pressure measure aspects of whole-person health, they do not provide a comprehensive assessment across various organ systems.
-The allostatic load index (ALI) is a composite measure of whole-person health, which reflects the multiple interrelated physiological regulatory systems that underlie healthy functioning [4, 5, 6].
-The ALI has been used to predict numerous health outcomes (e.g., [7, 8, 9, 10]) and measure racial disparities in health [11].
-Because it is sensitive to short-term interventions, the ALI is also an appealing “surrogate” outcome to test health interventions, like lifestyle and environmental modifications.
+The electronic frailty index (EFI) serves this purpose but is only designed for those aged 65 and older [5].
+A whole-person health metric for adults aged $18+$ would enable system-wide longitudinal tracking of individual and population health, comparisons across geographies and hospitals, and integration of an important predictor in algorithms relevant to the enterprise (e.g., predicting readmissions).
+Individual physiological variables measure aspects of health; they do not provide a comprehensive assessment across various organ systems.
+The allostatic load index (ALI) is a composite measure of whole-person health, which reflects the multiple interrelated physiological regulatory systems that underlie healthy functioning [6, 7, 8, 9].
+It has been used to predict numerous health outcomes (e.g., [10, 11, 12, 13]) and measure racial disparities in health [14].
+Because it is sensitive to short-term interventions, ALI is also an appealing “surrogate” outcome to test health interventions.
 Many ALI formulations exist.
-We focus on the calculation from Seeman et al.
-(2001) [5], which computes the ALI from ten component measurements across three body systems.
-As illustrated in Supplemental Figure S1, the original numeric measurements (e.g., hemoglobin A1c [HbA1c]) are discretized at clinically driven thresholds to define binary “stressor” indicators (e.g., $=1$ if hemoglobin A1c $\geq 6.5$ and $=0$ if hemoglobin A1c $<6.5$).
-The ALI is ultimately calculated as the sum of these ten stressors, i.e., the number of “unhealthy” measurements the patient had, with values of zero and ten representing the best and worst whole-person health, respectively.
+We focus on the calculation from [7], which computes ALI from ten components across three body systems (Table 1).
+The original numeric measurements (e.g., hemoglobin A1c [HbA1c]) are discretized at clinically-driven thresholds to define binary indicators of being at an unhealthy level (e.g., $=1$ if HbA1c $\geq 6.5$ and $=0$ if HbA1c $<6.5$).
+ALI is ultimately calculated as the sum of these ten components, with zero and ten representing the best and worst whole-person health, respectively (Supplemental Figure S1).
+Whole-person health is a latent concept influenced by genetics, environment, and other factors.
+Scores like the ALI incorporate components across body systems and yield “computable phenotypes” for it, concisely characterizing patients’ current health status with one-number summaries of observed data [15, 16].
+The ALI’s ability to quantify whole-person health has been rigorously examined in controlled studies [6, 7, 8].
+However, before adopting the ALI in other settings, it must be revalidated, accounting for real-world data challenges.
 1.2 Data Quality Challenges in Electronic Health Records Data
-EHR data present a huge opportunity to operationalize the ALI in the learning health system.
+EHR data present a substantial opportunity to operationalize the ALI in the learning health system.
 The volume of routinely collected data in the EHR is steadily climbing, increasing the accessibility to clinically meaningful variables.
-Analyzing data that come at little to no additional cost for collection is particularly appealing to biomedical researchers, leading to a surge in the utilization of EHR data in many clinical domains, including HIV and AIDS [12], genetics [13], emergency medicine [14], and therapeutic effectiveness [15].
+Analyzing data that come at little to no additional cost for collection is particularly appealing to biomedical researchers, leading to a surge in the utilization of EHR data in many clinical domains (e.g., [17, 18, 19, 20]).
 However, despite clear advantages, EHR data are collected in the fast-paced world of clinical care, leading to missingness and errors.
-These quality concerns remain a significant hurdle to the large-scale adoption of EHR in healthcare research [16, 17, 18, 19].
-In calculating the ALI, the original measurements may be (i) missing, making the stressor missing, or (i) non-missing but mismeasured, potentially misclassifying the stressor.
-Consider a patient whose true underlying HbA1c was $6.5\%$, such that their corresponding stressor should equal one.
+Quality concerns remain a significant hurdle to the large-scale adoption of EHR in healthcare research [21, 22, 23, 24].
+In calculating the ALI, measurements may be (i) missing, making the component missing, or (ii) non-missing but mismeasured, potentially misclassifying the component.
+Consider a patient whose true HbA1c was $6.5\%$, so their corresponding component should equal one.
 We illustrate two scenarios for this patient’s data in Figure 1.
-In the first, they had HbA1c measured, but the value was erroneously transposed to $5.6\%$.
-Based on the EHR, this patient’s stressor for HbA1c would be misclassified as zero since $5.6\%$ falls into the healthy range.
-In the second, they did not have HbA1c measured, so the value is missing, and we cannot define a stressor.
-Simply treating missing components as zero would also induce misclassification by treating them as “healthy." We accommodate missing data by instead defining the ALI as the proportion of non-missing stressors a patient experienced.
-Figure 1: Diagram of potential errors and missingness in a patient’s hemoglobin A1c (HbA1c) measurement in the electronic health record (EHR) before and after validation.
+In the first, HbA1c was measured but erroneously transposed to $5.6\%$.
+Based on the EHR, their HbA1c component would be misclassified as zero since $5.6\%$ falls into the healthy range.
+In the second, HbA1c was not measured, so the value is missing, and we cannot define a component.
+Counting missing components as zeros would also induce misclassification by treating them as “healthy." We accommodate missing data by instead defining ALI as the proportion of non-missing components a patient experienced.
+Figure 1: Diagram of potential errors and missingness in a patient’s hemoglobin A1c (HbA1c) measurement in the electronic health record (EHR) before and after validation (chart review).
 In the top scenario, the patient had HbA1c measured but with error, and the correct value was recovered through validation.
@@
 Errors in EHR data can also reduce the accuracy of prediction models [27].
-There is a particular concern for missingness and errors in derived variables, like the ALI, that combine data from many different variables [23, 24].
-Analyzing the ALI computed from the EHR data “as is” may misrepresent a patient’s actual health and lead to false conclusions about the association between ALI and healthcare utilization.
-For the ALI to be useful in research, it must be reliably and accurately measured.
-The original calculations, like Seeman et al.
-(2001) [5], were done in the context of clinical trials, where all patients must have all measurements and data validation protocols are likely in place to ensure quality.
-Similarly, validation, wherein the extracted EHR data are compared to original source documents (like the electronic chart) provides insights into data quality.
-Complete data validation (i.e., for all patients and variables) would ensure the integrity of EHR data but is overly expensive and unattainable in practice.
+There is a particular concern for missingness and errors in derived variables, like ALI, that combine data from many different variables [28, 29].
+Analyzing the ALI computed from the EHR data “as is” may misrepresent a patient’s actual health and lead to false conclusions about its association with healthcare utilization.
+Applying popular missing data approaches like multiple imputation (MI) may offer some improvement at baseline [30].
+However, they can also hinder results if any imputation models are misspecified [31].
+Since components are likely missing not at random in EHR data (e.g., clinicians usually order labs when they suspect values are abnormal) [32], correctly specifying imputation models could be especially challenging.
+Plus, MI would not correct for errors in non-missing ALI components.
+For the ALI to be useful in patient care and research, it must be reliably and accurately measured.
+The original calculations, like [7], were done in clinical trials, where all patients must have all measurements, and data validation protocols are likely in place to ensure quality.
+Similarly, validation via chart review, comparing extracted EHR data to source documents (like electronic charts), provides data quality insights.
+Complete data validation (for all patients and variables) would ensure the integrity of EHR data but is overly expensive and unattainable in practice.
 Fortunately, partially validating EHR data can help identify issues and pave the way for corrections.
 It provides a cost-effective alternative to complete data validation and is a promising approach to gauge the quality of observational data sources [33, 34, 35].
-These partial validation studies are a special case of two-phase designs [28, 29] (Figure 2A).
-Phase I includes variables available for all $N$ patients, which can be used to select a subsample of $n$ patients ($n<N$) for validation in Phase II.
-Figure 2: Diagram of two-phase designs with $N$ patients extracted from the electronic health records (EHR) in phase I and $n$ patients chosen for validation via chart review in phase II.
-The least informative strategy to select the validation study is A) simple random sampling.
-More informative targeted strategies include B) case-control sampling, C) balanced case-control sampling using discretized ALI, D) optimal stratified sampling using discretized ALI, and E) residual sampling based on the naive model.
-Within strata, the darker shaded areas indicate which patients would be selected.
+These partial validation studies are a special case of two-phase designs [36, 37].
+Phase I includes EHR data for all $N$ patients, which can be used to select a subsample of $n$ patients ($n<N$) for validation (chart reviews) in Phase II.
 Validated data are usually expected to be more correct (i.e., less error-prone) but not more complete (i.e., less missingness).
 To calculate the ALI from EHR data, we needed validation to do both.
-Herein, we devise a novel validation protocol that incorporates a “roadmap” created by clinical experts to recover missing data created by clinical experts using auxiliary health information (e.g., related diagnoses).
+We devise a novel validation protocol that incorporates a “roadmap” created by clinical experts to recover missing data using auxiliary information (e.g., related diagnoses).
+With it, we verify the ALI’s strength as a computable phenotype for whole-person health and a predictor of healthcare utilization in observational settings with data challenges.
 System Component Threshold Auxiliary Information
@@
 Vitamin deficiency
-Table 1: Ten component stressors of the allostatic load index were defined by discretizing measurements across three body systems at clinically driven thresholds.
-If the component was missing from the extracted EHR data, auditors searched for auxiliary information in the patient’s medical chart in Epic.
-If the auxiliary information was present, the component was treated as a “yes” and included in the patient’s validated ALI.
+Table 1: Ten components of the allostatic load index were defined by discretizing measurements across three body systems at clinically-driven thresholds.
+If the component was missing from the extracted EHR data, auditors searched for auxiliary information in the patient’s medical chart in Epic (the institution’s electronic charting software).
+If the auxiliary information was present, the patient was treated as having an “unhealthy” measurement and the component included in their validated ALI.
+1.3 Application to the Learning Health System
+We investigate how to overcome data challenges to measure whole-person health across one of the largest learning health systems in the United States.
+Having recently integrated the EFI into the EHR to measure whole-person health for patients $65$ and older [38], the health system is interested in operationalizing the ALI for the broader patient population.
+We obtained a representative sample of $N=1000$ patients from the EHR at Atrium Health Wake Forest Baptist Hospital in Winston-Salem, North Carolina.
+Our sample was drawn from patients (i) $18$-$65$ years old who (ii) initiated their first primary care outpatient encounter in the health system between March 11, 2018, and March 10, 2020 (to avoid data challenges and biases associated with the COVID-19 pandemic).
+Outpatient encounters were chosen to represent the typical patient voluntarily engaging in care.
+Our data contained demographics, vitals, lab values, discharge diagnoses, and problem lists dating $3$ years before the first encounter.
+Key variables included measurements for the ten ALI components (Table 1), age at first encounter (hereafter, “age”), and an indicator of having at least one hospitalization or ED visit during the $2$-year study period.
+Since most patients had multiple encounters, ALI components were based on mean measurements.
 1.4 Overview
-While very insightful, partial validation studies are expensive, time-intensive undertakings.
+While very insightful in the development and use of computable phenotypes, partial validation studies are expensive, time-intensive undertakings.
 Using a representative sample of $1000$ patients from the EHR at an extensive academic learning health system ($100$ of whom could be validated), we propose novel methods to design, conduct, and analyze statistically efficient and robust studies of whole-person health (ALI) and healthcare utilization.
-Using advanced statistics and bioinformatics methods, we seize three key opportunities to make the most of our investment: (i) robust and efficient statistical modeling, (ii) enriched validation protocol, and (iii) targeted validation study design.
+Using advanced statistics and bioinformatics methods, we seize three key opportunities to maximize our investment into partial validation: robust and efficient statistical modeling, enriched chart review protocol, and targeted sampling design.
 Statement of Significance
-Problem Standardized and valid summary measures of current health, i.e., whole-person health, are needed in the EHR to make patient- and institution-level decisions.
+  • Problem Standardized and valid summary measures of current health, i.e., whole-person health, are needed in electronic health records (EHR) to make patient- and institution-level decisions.
 However, routinely collected data are prone to missingness and errors.
@@
   • What This Paper Adds Building on the statistics and bioinformatics literature, we 1) devise a robust and efficient model for partially validated, error-prone EHR data; 2) propose a targeted validation study design to identify the most informative patients for our model; and 3) develop an enriched validation protocol that incorporates clinical expertise to improve the quality and completeness of EHR data.
+  • Who Would Benefit From the Knowledge in This Paper Biomedical researchers interested in analyzing routinely collected data, like those in the EHR, while correcting their data quality issues.
 2 Methods
-2.1 Model and Data
-This study was approved by the Institutional Review Board at Wake Forest University School of Medicine.
-We are interested in modeling healthcare utilization ($Y\in\{0,1\}$), defined as a binary indicator of having at least one hospitalization or emergency department (ED) visit in the $2$-year study period.
-The exposure of interest is the ALI ($X\in[0,1]$), defined as the proportion of stressors a patient experienced.
-Further, we control for patient age at first encounter ($Z\in\mathcal{R}$).
-These variables are assumed to be related through the usual logistic regression model for the log-odds that $Y=1$ (engaged in the healthcare system) given $X$ and $Z$.
-A sample of $N$ patients was extracted from the EHR to estimate these log odds ratios (ORs), denoted ${\boldsymbol{\beta}}$.
-Due to missingness and errors in the EHR, only an error-prone surrogate measure for $X$, denoted by $X^{*}$ ($X^{*}\in[0,1]$), is available.
-We compute $X^{*}$ as the proportion of non-missing stressors the patient experienced, according to the EHR data.
+2.1 Healthcare Utilization Model
+This study was approved by the Institutional Review Board at Wake Forest University School of Medicine (WFUSM).
+We model healthcare utilization ($Y\in\{0,1\}$), defined as a binary indicator of having at least one hospitalization or emergency department (ED) visit in the $2$-year study period.
+Whole-person health, measured by ALI ($X\in[0,1]$), is the exposure of interest.
+Each component $S_{j}$ ($j\in\{1,\dots,10\}$) equals one if the patient’s measurement was considered unhealthy and zero otherwise, such that $X=\sum_{j=1}^{10}S_{j}/10$ is the proportion of unhealthy measurements experienced.
+Further, we control for age ($Z\in\mathcal{R}$).
+We assume a logistic regression model for the log-odds that $Y=1$ (patient engaged in healthcare system) given ($X$, $Z$).
+A sample of $N$ patients was extracted from the EHR to estimate these log odds ratios (logORs), denoted ${\boldsymbol{\beta}}$.
 Age $Z$ is fully observed and expected to be error-free.
-To obtain consistent estimates for the desired healthcare utilization model, a two-phase design is planned (Figure 2A).
-In Phase I, the outcome $Y$ and surrogate exposure $X^{*}$ will be collected from the EHR for all $N$ patients.
-In Phase II, these data will be augmented with the validated exposure $X$ from a chart review for a subset of $n$ patients ($n<N$).
-To distinguish validated and unvalidated patients, let $V_{i}$ be a validation indicator, where $V_{i}=1$ if the $i$th patient is validated, $i\in\{1,\dots,N\}$, and $V_{i}=0$ otherwise.
-Incorporating data from Phases I and II, along with the validation indicator, the joint distribution of a complete observation is assumed to be
+However, due to missingness and errors in ALI components, only an error-prone measure for $X$, denoted by $X^{*}$ ($X^{*}\in[0,1]$), is available from the EHR data.
+We define error-prone components $S_{j}^{*}$ ($j\in\{1,\dots,10\}$) and introduce missingness indicators $M_{j}^{*}$ for them (i.e., $M_{j}^{*}=1$ if $S_{j}^{*}$ is missing).
+We compute $X^{*}=\sum_{j=1}^{10}(1-M_{j}^{*})S_{j}^{*}/\sum_{j=1}^{10}(1-M_{j}^{*})$, i.e., as the proportion of non-missing components experienced.
+Calculating ALI as a proportion – rather than sum – was critical.
+Summing over non-missing components treats missing values as zeros (i.e., healthy measurements); instead, the proportion ignores them.
+2.2 Three Key Steps to Partial Validation Studies: Analysis, Design, and Protocol
+To obtain consistent estimates for the healthcare utilization model, we planned a partial validation study.
+We first extracted EHR data ($Y$, $X^{*}$, $Z$) for a sample of $N=1000$ patients.
+Then, a subset of $n=100$ patients was selected for validation via chart review such that $X$ was also measured.
+Partial validation studies break down into three key steps.
+Design: Select which patients should be validated.
+Protocol: Develop standardized tools and practices for those patients’ chart reviews.
+Analysis: Estimate statistical models or quantities of interest from validated and unvalidated data together.
+These steps are intertwined, with design choices depending on the planned analysis, and each step provides opportunities to gain information.
+For added flexibility, a multi-wave validation framework, conducting chart reviews in multiple sprints, allows us to revisit key steps and make modifications as information accumulates (Figure 2).
+In our application (Section 3), we built the statistical model and developed the baseline sampling strategy and tools before beginning validation.
+The proposed methods are detailed in order of development: analysis (Section 2.3), design (Section 2.4), and protocol (Section 2.5).
+Across three validation waves, we updated our sampling strategy and tools (revisiting the design and protocol steps, respectively), before fitting the final model (Section 2.6).
+Figure 2: After extracting the electronic health records (EHR) data, partial validation studies break down into three key steps: analysis, design, and protocol.
+These steps are intertwined, with design choices depending on the planned analysis, and each step provides opportunities to gain information.
+For added flexibility, a multi-wave validation framework allows us to revisit these key steps and make modifications as information accumulates.
+2.3 Analysis: Robust and Efficient Statistical Model for Partially Validated Data
+Let $V_{i}$ distinguish validated and unvalidated patients, where $V_{i}=1$ if the $i$th patient is validated, $i\in\{1,\dots,N\}$, and $V_{i}=0$ otherwise.
+Incorporating EHR and chart review data, along with the validation indicator, the joint distribution of a complete observation is assumed to be
 $\Pr(V,X^{*},Y,X,Z)$ $=\Pr(V\mid Y,X^{*},Z){\Pr}_{{\boldsymbol{\beta}}}(Y\mid X,Z)\Pr(X\mid X^{*},Z)\Pr(X^{*},Z),$ (1)
-where $\Pr(V\mid Y,X^{*},Z)$ is the validation sampling probability; ${\Pr}_{{\boldsymbol{\beta}}}(Y\mid X,Z)$ corresponds to the logistic regression model; $\Pr(X\mid X^{*},Z)$ is the exposure error mechanism; and $\Pr(X^{*},Z)$ is the joint probability density or mass function of $X^{*}$ and $Z$.
-Two assumptions are made in defining (1).
-First, selection for validation (i.e., $V$) is assumed to depend only on Phase I variables ($Y$, $X^{*}$, $Z$), so the validated exposure $X$ is missing at random (MAR) for unvalidated patients [30].
-Second, the exposure $X^{*}$ from the EHR is assumed to be a surrogate for the validated exposure $X$, such that $Y$ and $X^{*}$ are conditionally independent given $X$.
-In other words, once $X$ is known, $X^{*}$ is assumed not to contribute anything further about the relationship with $Y$.
-While other ways of factoring (1) are possible, this factorization avoids making an additional assumption about $\Pr(X^{*},Z)$.
-2.2 Robust Statistical Model for Partially Validated Data
-The observed-data log-likelihood function (hereafter, log-likelihood function) for the log ORs ${\boldsymbol{\beta}}$ is defined based on the joint distribution in (1), but first, two simplifications can be made.
-First, the validation sampling probability $\Pr(V|Y,X^{*},Z)$ can be omitted from the log-likelihood for ${\boldsymbol{\beta}}$ because the Phase II variables are MAR.
-Second, the joint probability $\Pr(X^{*},Z)$ can be ignored because it will drop out as the log-likelihood function is maximized with respect to ${\boldsymbol{\beta}}$ since $X^{*}$ and $Z$ are fully observed.
-Observations for all $N$ patients in the study are assumed to be independent and identically distributed (i.i.d.) following (1).
-Now, the log-likelihood function for ${\boldsymbol{\beta}}$ is proportional to the following:
+where $\Pr(V\mid Y,X^{*},Z)$ is the validation sampling probability; ${\Pr}_{{\boldsymbol{\beta}}}(Y\mid X,Z)$ is the logistic regression model; $\Pr(X\mid X^{*},Z)$ is the exposure error mechanism; and $\Pr(X^{*},Z)$ is the joint probability density or mass function of ($X^{*}$, $Z$).
+Two assumptions are made in (1).
+First, selection for validation depends only on fully observed variables ($Y$, $X^{*}$, $Z$), so the validated exposure $X$ is missing at random (MAR) for unvalidated patients [39].
+Second, $X^{*}$ from the EHR is a surrogate for validated $X$ from the chart reviews, such that $Y$ and $X^{*}$ are conditionally independent given $X$.
+Observations for all $N=1000$ patients are assumed to be independent and identically distributed following (1).
+Thus, the observed-data log-likelihood function (hereafter, log-likelihood) for the logORs ${\boldsymbol{\beta}}$ is defined from (1), after two simplifications.
+First, we can omit $\Pr(V|Y,X^{*},Z)$ from the log-likelihood because $X$ is MAR.
+Second, we can ignore $\Pr(X^{*},Z)$ because it drops out as the log-likelihood is maximized with respect to ${\boldsymbol{\beta}}$ since ($X^{*}$, $Z$) are fully observed.
+Now, the log-likelihood for ${\boldsymbol{\beta}}$ is proportional to
 $\sum_{i=1}^{N}V_{i}\left[\log\left\{{\Pr}_{{\boldsymbol{\beta}}}(Y_{i}\mid X_{i},Z_{i})\right\}+\log\left\{\Pr(X_{i}\mid X_{i}^{*},Z_{i})\right\}\right]$
 $+\sum_{i=1}^{N}(1-V_{i})\log\left\{\sum_{x=0}^{1}{\Pr}_{{\boldsymbol{\beta}}}(Y_{i}\mid x,Z_{i})\Pr(x\mid X_{i}^{*},Z_{i})\right\}.$ (2)
-Notice that unvalidated patients contribute their incomplete information to this log-likelihood by summing over the missing validated exposure $X$:
+Notice that unvalidated patients contribute incomplete information to (2) by summing over the missing validated $X$:
 $\sum_{x=0}^{1}{\Pr}_{{\boldsymbol{\beta}}}(Y_{i}\mid x,Z_{i})\Pr(x\mid X_{i}^{*},Z_{i})$ $\propto\sum_{x=0}^{1}{\Pr}_{{\boldsymbol{\beta}}}(Y_{i},x,Z_{i},X_{i}^{*})={\Pr}_{{\boldsymbol{\beta}}}(Y_{i},Z_{i},X_{i}^{*}),$ (3)
 remembering that $X$ takes on values $x\in\{0,0.1,\dots,0.9,1\}$ here.
-Still, a model for the exposure error mechanism $\Pr(X\mid X^{*},Z)$ is needed to estimate the analysis model parameters since $X$ is only observed for the subset of validated patients.
-Since the exposure $X$, representing the patient’s ALI, is bounded between zero and one (inclusive), there are not many flexible, parametric choices to model the distribution of $X$ given ($X^{*},Z$).
-Thus, a nonparametric estimator is preferable, allowing the validated exposure $X$ to take on any values and arbitrarily depend on $X^{*}$ and $Z$.
-Lotspeich et al.
-(2022) [31] proposed a semiparametric approach to logistic regression with outcome and exposure errors, which nonparametrically (i.e., without assuming a particular distribution or structure) estimates the exposure error mechanism as described.
+A model for $\Pr(X\mid X^{*},Z)$ is needed to estimate ${\boldsymbol{\beta}}$ from (2) since $X$ is only observed for validated patients.
+Since $X\in[0,1]$, there are not many flexible, parametric choices to model its distribution given ($X^{*},Z$).
+A nonparametric estimator is preferable, allowing $X$ to take on any values and arbitrarily depend on ($X^{*}$, $Z$).
+[40] proposed a semiparametric approach to logistic regression with outcome and exposure errors, which nonparametrically (i.e., without assuming a particular distribution or structure) estimates the exposure error mechanism as described.
 The current setting is a special case with only exposure error, and a brief outline is provided in the Supplemental Materials.
-For full technical details on the semiparametric sieve maximum likelihood estimators (SMLEs), see Lotspeich et al.
-(2022) [31].
-2.3 Targeted Designs to Select Informative Patients for Validation
-As a baseline, consider the simple random sampling (SRS) design, wherein $n$ of the $N$ patients are chosen for validation with equal probability.
-This strategy ignores all available information, making it the most uninformative.
-For full-likelihood approaches like the SMLE, any information available for all patients can – and should – be used to design the validation study.
+The logiSieve package contains R functions to fit this model [41].
+For full technical details on the semiparametric sieve maximum likelihood estimators (SMLEs), see [40].
+2.4 Design: Targeted Sampling to Select Informative Patients for Validation
+Based on our analysis plan, we considered various sampling designs to select the most informative $100$ patients for validation.
+When validating computable phenotypes, simple random sampling (SRS) is often used.
+Under SRS, $100$ patients would be chosen for chart review with equal probability (Figure 3A).
+This strategy ignores all available information.
+Some phenotyping algorithms require an SRS validation study.
+However, for full-likelihood approaches like SMLEs, any information available for all patients can be used to design the validation study.
 Designing based on fully observed information ensures that the MAR assumption holds for the downstream analysis while allowing for many targeted strategies.
+Figure 3: Diagram of partial validation studies with $N$ patients extracted from the electronic health records (EHR) and $n$ patients chosen for validation via chart review.
+The least informative strategy to select the validation study is A) simple random sampling.
+More informative targeted strategies include B) case-control sampling, C) balanced case-control sampling using discretized ALI, D) optimal stratified sampling using discretized ALI, and E) residual sampling based on the naive model.
+Within strata, the darker shaded areas indicate which patients would be selected.
 When designing a validation study to promote statistical efficiency of the parameter estimates (i.e., to minimize the variance in the estimated logOR on ALI), we start at the end.
-We begin with the planned logistic regression model of $Y$ given $X$ and $Z$, and work backward to determine which patients will be most informative to it.
-For this reason, the “best” designs vary across different estimators, with model-based estimators, like the SMLEs, offering the most design flexibility.
-Other estimators that rely on the design–rather than a model–to handle the missingness, like inverse probability weighting, require nonzero sampling probability for all patients, but model-based estimators do not.
-We considered stratified and rank-based validation study designs.
+Beginning with the planned logistic regression model of $Y$ given ($X$, $Z$), we work backward to determine which patients will be most informative.
+For this reason, the “best” designs vary across different estimators, with model-based ones, like SMLEs, offering the most design flexibility.
+Estimators that rely on the design to handle missingness, like inverse probability weighting (IPW), require nonzero sampling probability for all patients; model-based estimators do not.
+We considered stratified and rank-based designs.
 Multiple validation waves were conducted to allow the protocol and study design to adapt as information accumulated, which is becoming common in this space [42, 34, 43].
-Ultimately, simulations based on preliminary validation data (see Section 3.3.3) guided our selection of the targeted design used in the application.
-Given our binary outcome, the case-control (CC) design is a natural option here, wherein $n/2$ patients each are chosen from subgroups with $Y=0$ and $Y=1$ (Figure 2B).
-The balanced case-control (BCC) design further stratifies on covariate information, subsequently choosing equal (i.e., balanced) numbers from each stratum (Figure 2C).
-When the covariates are continuous, they must be discretized to form strata.
+Ultimately, simulations based on preliminary validation data guided our selection of the targeted design used in the application (see Section 2.6).
+2.4.1 Stratified Designs
+Given our binary outcome, the case-control (CC) design is a natural option, wherein $50$ patients each are chosen from subgroups with $Y=0$ and $Y=1$ (Figure 3B).
+The balanced case-control (BCC) design further stratifies on covariates, subsequently choosing equal (i.e., balanced) numbers from each stratum (Figure 3C).
+Continuous covariates must be discretized to form strata.
 For instance, we could use $X_{D}^{*}$, a discretized version of $X^{*}$, to define strata of patients with error-prone ALI above and below the average.
-Then, $n/4$ patients are validated from each of the four $(Y,X_{D}^{*})$ strata.
-Importantly, we can still fit the model using the original continuous values.
-While the CC and BCC designs are intuitive and can offer some efficiency gains over SRS, there are slightly more complicated designs that are guaranteed to be theoretically “optimal,” meaning that they choose the number of patients to validate from each stratum to achieve the lowest possible variance for the log OR on ALI.
-Many have been proposed to select patients for the collection of “expensive” binary exposures, like biomarker measurement or validation via chart review, for a variety of analytical approaches to modeling binary outcomes [34, 35, 36, 37, 29, 38].
-Recently, Lotspeich et al.
-(2024) [33] proposed a design for optimal validation of secondary use data for maximum likelihood estimation with binary outcome and exposure misclassification (Figure 2D), which could be applied here.
-However, we would achieve the lowest possible variance for the discretized version $X_{D}^{*}$ rather than $X^{*}$.
-Increasing the number of unique values of $X_{D}^{*}$ (i.e., defining more strata) can help, but information will still be lost.
-Still, designs that rank patients rather than stratifying them offer improved granularity and sizable efficiency gains.
-For example, extreme tail sampling (ETS) designs focus on the extremes (i.e., tails) of the distribution, ranking the data by some value of interest and targeting patients with the smallest or largest values.
-To identify patients who are most informative not just about ALI ($X^{*}$ and $X$) but about its relationship to healthcare utilization ($Y$), residual sampling (RS) is preferred [39].
-With measurement error, we begin by fitting the “naive” analysis model for $Y$ (i.e., using $X^{*}$ instead of $X$) to obtain estimates for the log ORs $\hat{\boldsymbol{\beta}}^{*}$.
-Then, residuals for the naive model are computed for each patient $i$ ($i=1,\dots,N$) as
+Then, validate $25$ patients from each of the four ($Y$, $X_{D}^{*}$) strata.
+Importantly, we can still fit the model using the continuous values $X^{*}$.
+The CC and BCC designs are intuitive and offer efficiency gains over SRS.
+Still, slightly more complicated designs are guaranteed to be theoretically “optimal;” they choose the number of patients from each stratum to achieve the lowest possible variance for the logOR on ALI [44, 45, 46, 47, 37, 48].
+[43] proposed the optimal (OPT) design for model-based analysis of data with binary outcome and exposure misclassification (Figure 3D), which could apply here.
+However, we would achieve the lowest possible variance for $X_{D}^{*}$ rather than $X^{*}$.
+Increasing the number of unique values of $X_{D}^{*}$ (i.e., more strata) can help, but information will still be lost.
+Moreover, the OPT design depends on the true logORs, which are unknown; reliable placeholders for them are required to implement the design (often collected in a multi-wave framework).
+2.4.2 Rank-Based Designs
+Designs that rank patients rather than stratifying them offer improved granularity and sizable efficiency gains.
+For example, extreme tail sampling (ETS) designs rank patients by some value of interest and target those with the smallest or largest (i.e., most “extreme”) values.
+To identify patients most informative not just about ALI but its relationship to healthcare utilization, residual sampling (RS) is preferred [49].
+To implement RS with measurement error, we begin by fitting the “naive” analysis model, i.e., using standard logistic regression procedures with $Y$, $X^{*}$ (in place of $X$), and $Z$ to obtain estimates for the logORs $\hat{\boldsymbol{\beta}}^{*}$ from the full sample of $1000$ patients.
+Then, the naive model residuals are computed for each patient $i$ as
 $r_{i}^{*}$ $=Y_{i}-{\Pr}_{\hat{\boldsymbol{\beta}}^{*}}(Y_{i}=1|X^{*}_{i},Z_{i})$
-$=Y_{i}-\frac{1}{1+\exp\left\{-\left(\beta_{0}^{*}+\beta_{1}^{*}X_{i}^{*}+{\boldsymbol{\beta}}_{2}^{*}Z_{i}\right)\right\}},$ (4)
-and used to rank the $N$ patients in the study.
-With logistic regression, the largest residuals $r_{i}^{*}$ (in magnitude) correspond to observations where the observed value $Y_{i}$ had a low predicted probability (Figure 2E).
-The auditDesignR package, available on GitHub at www.github.com/sarahlotspeich/auditDesignR, contains R functions to sample from all of these designs.
-2.4 Enriched Validation Protocol to Reduce Errors and Missingness
-The chart reviews to validate the EHR data were conducted over $6$ months (February to July 2024) by a team of four clinical research assistants and technicians in the Wake Forest University School of Medicine (hereafter, “auditors”).
-These skilled auditors were chosen for their familiarity with Epic, the institution’s electronic charting software, and their demonstrated experience in the clinical research environment.
+$=Y_{i}-\left[1+\exp\left\{-\left(\hat{\beta}_{0}^{*}+\hat{\beta}_{1}^{*}X_{i}^{*}+\hat{\beta}_{2}^{*}Z_{i}\right)\right\}\right]^{-1},$ (4)
+and used to rank patients.
+The largest residuals $r_{i}^{*}$ (in magnitude) correspond to patients whose observed outcome $Y_{i}$ had a low predicted probability (Figure 3E).
+The auditDesignR package contains R functions to sample from all of these designs [50].
+2.5 Protocol: Enriched Validation Protocol to Reduce Errors and Missingness
+Chart reviews were conducted over $6$ months by four clinical research assistants and technicians at WFUSM (hereafter, “auditors”), chosen for their familiarity with Epic (the institution’s electronic charting software) and demonstrated research experience.
 2.5.1 Identifying Sources of Auxiliary Information
 Before the chart reviews, our clinical collaborators developed the “audit roadmap” (Table 1).
-While it is unlikely that we would identify the original missing measurements, we harnessed the clinical expertise of our team to identify sources of auxiliary information about the missing discretized components.
+While it is unlikely that we would identify the original missing measurements, we harnessed clinical expertise to identify auxiliary information about missing discretized components.
 The roadmap suggested where auditors could look in Epic to locate supplemental information about missing values.
-For example, if a patient were missing HbA1c, the auditors would search their chart for a diagnosis of diabetes or impaired glycemic control.
-If either were present, the patient’s diagnosis would suggest their missing HbA1c value would have exceeded the clinically meaningful threshold of $6.5\%$.
+It employs “anchors” to try and recover missing components using clinically-driven placeholders [15].
+For example, if a patient were missing HbA1c, auditors would search their chart for a diagnosis of diabetes or impaired glycemic control.
+If either was present, the patient’s diagnosis would suggest that their missing HbA1c value would have exceeded the $6.5\%$ threshold.
 Thus, their validated ALI $X$ would include one point for HbA1c.
-If both were absent, the patient’s validated ALI would continue to exclude this still-missing component.
-Many components, like C-reactive protein and HbA1c, had multiple sources of auxiliary information, while others, like triglycerides and total cholesterol, had just one.
+If neither diagnosis was present, $X$ would continue to exclude this still-missing component.
+Many components, like HbA1c, had multiple sources of auxiliary information, while others, like triglycerides, had just one.
 Only serum albumin, a blood test to measure liver and kidney function, did not have a supplemental diagnosis.
 2.5.2 Validated Data Collection
-The EHR data were extracted by the Wake Forest Clinical and Translational Science Institute and securely received by the analytical team.
-In preparation for validation, the analytical team transformed the data from the original wide format (with one row per patient encounter and columns for different variables) to a long format (with one row per patient encounter per variable).
-This transformation streamlined the data structure for the auditors, with only one value to validate per row (Figure 3).
-Figure 3: Illustrative example of how the original extracted EHR data for patients’ vitals were transformed from a wide format (with one row per patient encounter and columns for different variables) to a long format (with one row per patient encounter per variable), merged with the audit roadmap, and augmented with empty columns for auditors to enter their findings.
-The same was done with patients’ lab data.
-Before sending the data to the auditors, the roadmap was incorporated, and an empty column was added for the “reviewed value” (i.e., the one found in Epic).
-Depending on the scenario, the auditors would enter different information for the reviewed value, and each data point was assigned an audit finding (flow chart in Figure 4).
-Three findings were possible for non-missing components in the EHR.
-  1. 1.
-    Extracted value correct: If the component matched Epic, the auditor would enter the original value again.
-  2. 2.
-    Extracted value incorrect: If the component did not match Epic, the auditor would enter the new value found in Epic.
-  3. 3.
-    Extracted value not found: If the component could not be found in Epic, the auditor would enter “Not Found.”
-Two findings were possible for missing components.
-  1. 4.
-    Auxiliary information found: If auxiliary information from the audit roadmap were found in Epic, the auditor would enter the auxiliary information.
-  2. 5.
-    Auxiliary information not found: If auxiliary information from the audit roadmap were not found in Epic, the auditor would enter “Not Found.”
-Figure 4: Flow chart of instructions given to auditors for the chart reviews (Steps 1–3) and the corresponding audit finding assigned by the analysts (Step 4). Column names refer to those in Figure 3.
-3 Results
-3.1 Cohort Description
-We obtained a representative sample of $N=1000$ patients from the EHR at Atrium Health Wake Forest Baptist Hospital in Winston-Salem, North Carolina.
-To be considered, patients needed to (i) be $18-65$ years old and (ii) initiate their first primary care outpatient encounter in the health system between March 11, 2018, and March 10, 2020 (to avoid data challenges and biases associated with the COVID-19 pandemic).
-Outpatient encounters were chosen to represent the typical patient voluntarily engaging in care.
-Our data contained demographics, vitals, lab values, discharge diagnoses, and problem list values dating $3$ years before the first encounter.
-Key variables included the measurements for the ten components of the ALI (Table 1), age at first encounter (hereafter, “age”), and an indicator of having at least one hospitalization or ED visit during the $2$-year study period.
-Since most patients had multiple encounters, the ALI components were based on mean measurements.
-Sampled patients had a median age of $48$ years and a median income of \$52,320.
-They were $61\%$ female, $72\%$ White, $6\%$ Hispanic, and $73\%$ homeowners.
-There were $318$ patients ($32\%$) with at least one hospitalization or ED visit.
-Based on the original unvalidated EHR data, the distribution of error-prone ALI was right-skewed (Supplemental Figure S2).
-With a median of $0.33$ (interquartile range $=0.17,0.50$), on average, patients in our study experienced $1$ out of every $3$ non-missing stressors on their whole-person health.
-3.2 Baseline Data Quality and Naive Model Estimates
-As expected, the extracted EHR data had a large amount of missingness.
-In particular, rare labs for inflammation components were missing for $>95\%$ of patients.
-Labs for the metabolic components were missing for $21\%$ of patients, and measurements for the cardiovascular components were non-missing for basically everyone (Supplemental Figure S3A).
-The most common patterns of missing data are displayed in Supplemental Figure S4.
-Fortunately, most patients ($78\%$) had at least $5$ non-missing ALI components (Supplemental Figure S3B).
-Using these ALI values to fit the “naive” healthcare utilization model, we estimate the following.
-First, for every $0.1$-point (or one component) increase in ALI, the odds of engaging in the healthcare system are expected to increase by $10\%$ (OR $=1.10$, 95% confidence interval [95% CI]: $1.03$, $1.18$), adjusting for age.
-Second, the expected odds of engaging in the healthcare system for an $18$-year-old patient with an ALI of 00 are $0.25$ (95% CI: $0.18$, $0.35$).
-Third, for every $10$ years older a patient was at their first encounter, their odds of engaging in the healthcare system are expected to increase by $11\%$ (OR $=1.11$, 95% CI: $1.00,1.23$), adjusting for ALI.
-Thus, the naive model suggests that worse whole-person health leads to higher expected odds of hospitalization or emergency department visits.
-3.3 Three Waves of Data Validation Via Chart Review
-While the naive model aligns with the existing literature, we wanted to ensure that the missingness and potential errors in the EHR data did not lead to spurious conclusions.
-Thus, we secured funding to validate data for $n=100$ of the $N=1000$ patients, whose selection and review were broken down into three waves: (i) Pilot, (ii) Wave I, and (iii) Wave II.
-3.3.1 Pilot Wave: Testing the Audit Protocol and Establishing Inter-Auditor Agreement
-The first $n_{0}=4$ patients were selected for chart review via the BCC design introduced in Section 2.3.
-Specifically, patients were split into four strata based on their outcome and error-prone ALI, discretized at the median:
-  • (i)
-    Engaged in the healthcare system and low error-prone ALI ($Y=1$, $X^{*}\leq 0.33$),
-  • (ii)
-    Engaged in the healthcare system and high error-prone ALI ($Y=1$, $X^{*}>0.33$),
-  • (iii)
-    Did not engage in the healthcare system and low error-prone ALI ($Y=0$, $X^{*}\leq 0.33$), and
-  • (iv)
-    Did not engage in the healthcare system and high error-prone ALI ($Y=0$, $X^{*}>0.33$).
-Then, $1$ patient was randomly chosen for validation from each stratum.
-The auditors followed the protocol in Section 2.4 to (i) verify the non-missing values in the EHR data or note any discrepancies and (ii) locate auxiliary information for missing values.
-Following the completion of the Pilot wave, the auditors and analysts met to discuss any modifications that would be helpful before proceeding.
-For example, the third audit finding of “extracted value not found” was not one of the original options, and the team discussed how these data points should be recorded.
-Adding an optional “Notes” column, where auditors could include any additional information they uncovered, also resulted from this discussion.
-The Pilot and Wave I findings were combined to interpret data quality (see Section 3.3.2).
+In preparation for validation, the analytical team transformed the extracted EHR data from the original wide format (one row per patient encounter and columns for different variables) to a long format (one row per patient encounter per variable).
+This transformation streamlined the data structure for auditors: one value to review per row (Supplemental Figure S2).
+Before sending the data to the auditors, the roadmap was incorporated, and an empty column was added for the “reviewed value” (i.e., the one in Epic).
+Depending on the scenario, auditors would enter different information for the reviewed value, and each data point was assigned one of five auditor findings (Figure 4).
+Figure 4: Flow chart of instructions given to auditors for the chart reviews (Steps 1–3) and the corresponding auditor finding assigned by the analysts (Step 4).
+Column names refer to those in Supplemental Figure S2.
+2.6 Multi-Wave Strategy: Sampling Adaptively and Improving Tools
+Choosing the “best” design often relies on information that is unknown at baseline, like EHR error rates or the magnitude of the association between ALI and healthcare utilization.
+Since we did not have any prior information about these unknowns, we started by sampling conservatively to collect preliminary information for a more extreme, targeted design later on.
+Multi-wave validation schemes like this one are appealing because they balance learning about our data with strategically designing for them [42, 34].
+Even if prior information is available (e.g., from previous chart reviews), multi-wave validation allows us to relearn how data have changed before dedicating resources to validate them [43].
+Further, we can seize opportunities to incorporate auditors’ feedback and improve the validation protocol.
+We validated $100$ patients’ data across three waves: the Pilot, Wave I, and Wave II.
+The small Pilot wave tested the validation protocol and established inter-auditor agreement on the first four patients.
+Then, Waves I and II put this protocol into action, as auditors validated the remaining patients.
+Waves I and II each selected $48$ more patients.
+(Designating equal-sized waves balances design flexibility and stability [42].) In Wave II, information from the Pilot + Wave I was leveraged to design a more targeted strategy.
+Patients could be chosen for validation at most once.
+2.6.1 Adapting the Validation Study Design as Information Accumulates
+Between the Pilot and Wave I, the first $52$ patients were selected for chart review via the BCC design.
+Specifically, patients were split into four strata based on their outcome and error-prone ALI (discretized at the median of $0.33$).
+Then, equal numbers of patients were randomly chosen from the four strata.
+In the Pilot, one patient was chosen from each stratum; another $12$ were chosen per-stratum in Wave I.
+Before validating the remaining $48$ patients, we designed bespoke simulation studies based on the preliminary validation data.
+They explored how designs compared under different scenarios for true underlying data quality.
+The design offering the best efficiency (i.e., lowest variability) in estimating the logOR on ALI with the SMLE was desired.
+Under varied error rates, the targeted validation study designs offered empirically unbiased estimates and higher efficiency than SRS.
+The RS design provided the best efficiency under extra low to moderate error settings, and the preliminary validation data (Pilot + Wave I) indicated a perfect true positive rate (TPR) $=1.00$ and negligible false positive rate (FPR) $<0.01$.
+Under full/high data recovery rates ($90$-$100\%$ missing data recovered through validation) and different designs, the SMLE was empirically unbiased.
+While biased for moderate recovery or lower ($\leq 50\%$), the SMLE was still closer to the truth than the naive estimate was (i.e., based on unvalidated EHR data).
+Regardless of how much missing data was recovered, the RS design offered the highest efficiency.
+Details like the data-generating mechanisms can be found in the Supplemental Materials.
+We concluded that the RS design was the most informative way to select the remaining $48$ patients for validation in Wave II.
+To implement it, each unvalidated patient’s residual was calculated based on the naive healthcare utilization model estimates (Section 3.3) following (4), as
+$r_{i}^{*}=Y_{i}-\frac{1}{1+\exp\left\{-\left(-1.383+0.945X_{i}^{*}+0.103Z_{i}\right)\right\}}.$
+Then, patients were ranked by $r_{i}^{*}$, and the $24$ patients with the smallest residuals and the $24$ patients with the largest residuals were selected for validation.
+Here, the RS design prioritized (i) patients with zero hospitalizations or ED visits but high predicted probabilities of having at least one and (ii) patients with at least one hospitalization or ED visit but high predicted probabilities of having zero.
+That is, patients whose observed $Y_{i}$ aligned well with the naive model’s prediction were not considered to be very informative and thus were deprioritized for validation.
+2.6.2 Testing the Validation Protocol and Establishing Inter-Auditor Agreement
+The validation protocol evolved slightly throughout the waves, as well.
+After completing the first four patients’ chart reviews (Pilot), auditors and analysts met to discuss any protocol modifications that would be helpful before proceeding.
+For example, the auditor finding “extracted value not found” was not one of the original options, and the team discussed how these data points should be recorded.
+Adding an optional “Notes” column, where auditors could include any additional information, also resulted from this discussion.
+The Pilot also presented a valuable opportunity to evaluate inter-auditor agreement.
 We started with two auditors on our team.
@@
 Ultimately, agreement in the four auditors’ findings was calculated using Fleiss’ Kappa [51], as detailed in the Supplemental Materials.
-We found high agreement between the auditors’ findings for lab ($\kappa=0.90$) and vitals ($\kappa=0.81$) variables.
-3.3.2 Wave I: Collecting Preliminary Audit Data
-The next $n_{1}=48$ patients validated were also selected via the BCC design from Section 3.3.1.
-Since we did not have any prior information about the quality of the EHR data, we sampled more conservatively here to collect preliminary information for a more extreme, targeted design in Wave II.
-Multi-wave validation schemes like this one are appealing because they balance learning about our data with strategically designing for them [32, 26].
-Even if prior information is available (e.g., from a previous audit), validating in multiple waves allows us to relearn how our data may have changed before using all our resources to validate them [33].
+We found high agreement between the auditors’ findings for labs ($\kappa=0.90$) and vitals ($\kappa=0.81$).
+3 Results
+3.1 Cohort Description and Study Objectives
+We considered adult patients ($18$-$65$ years old) with at least one primary care outpatient encounter at Atrium Health Wake Forest Baptist Hospital.
+The $1000$ sampled patients had a median age of $48$ years and a median income of \$52,320.
+They were $61\%$ female, $72\%$ White, and $6\%$ Hispanic.
+During the $2$-year study period, $318$ patients ($32\%$) had at least one hospitalization or ED visit.
+In unvalidated EHR data, error-prone ALI was right-skewed (Supplemental Figure S3).
+With a median of $0.33$ (interquartile range $=0.17,0.50$), patients experienced one out of every three non-missing components on their whole-person health, on average.
+We validated ALI data for $100$ patients.
+First, we discuss the EHR data quality (Section 3.2).
+Second, we present the healthcare utilization model (Section 3.3).
+Within these sections, results are discussed chronologically, beginning with findings from the unvalidated EHR data and ending with those following the completion of all validation waves.
+We also discuss how these findings inform future implementations of the ALI in this healthcare system (Section 3.4).
+3.2 Validation Study Findings: EHR Data Quality
+As expected, the extracted EHR data had lots of missingness.
+Rare labs for inflammation components were missing for $>95\%$ of patients (Supplemental Figure S4A).
+Labs for metabolic components were missing for $21\%$ of patients.
+Common patterns of missing data are displayed in Supplemental Figure S5.
+Measurements for the cardiovascular components were non-missing for basically everyone.
+Fortunately, most patients ($78\%$) had at least five non-missing ALI components (Supplemental Figure S4B).
+3.2.1 Pilot + Wave I Chart Reviews
 […]
@@
 These findings refer to the original numeric measurements (before discretizing them into allostatic load index components), and there could be multiple per patient per variable.
-Fewer vitals than labs were found to match Epic ($59\%$ versus $81\%$), with $31\%$ not found.
-Through validation, we discovered that the EHR data extraction algorithm expected every patient encounter to have vitals.
-For encounters where vitals were not taken (e.g., physical therapy or optometry visits), values were imputed from other encounters; these imputed values could not be found in Epic by the auditors.
+Auditor findings were mutually exclusive and specific to whether the measurements were originally non-missing (“extracted value correct,” “extracted value incorrect,” or “extracted value not found”) or missing (“no auxiliary information found” or “auxiliary information found”) in the extracted EHR data.
+Figure 6: Comparison of the discretized allostatic load index components from the unvalidated EHR data to A) $52$ patients’ data in the Pilot + Wave I, B) $48$ patients’ data in Wave II, and C) $100$ patients’ data in all waves combined.
+Fewer vitals than labs matched Epic ($59\%$ versus $81\%$), with $31\%$ not found.
+We discovered that the EHR data extraction algorithm expected every patient encounter to have vitals.
+For encounters where vitals were not taken (e.g., physical therapy), values were imputed from other encounters; imputed values could not be found in Epic by the auditors.
 Thus, we uncovered an area for potential improvement in the current EHR data extraction protocol.
-Figure 6: Comparison of the discretized allostatic load index components from the unvalidated EHR data to A) $52$ patients’ data in the Pilot + Wave I, B) $48$ patients’ data in Wave II, and C) $100$ patients’ data in all waves combined.
-Incorporating the audit roadmap into the validation protocol improved the completeness of the data, as the number of missing values decreased for nearly all components of the ALI (Supplemental Figure S3A).
-Homocysteine saw the biggest improvement, with auxiliary information located for $23$ of the $52$ validated patients ($44\%$).
+Incorporating the audit roadmap into the validation protocol improved the completeness of the data.
+The number of missing values decreased for nearly all ALI components (Supplemental Figure S4A).
+Homocysteine saw the biggest improvement, with auxiliary information located for $23$ of $52$ validated patients ($44\%$).
 The number of non-missing ALI components per patient increased from $6.6$ to $7.5$, on average.
 Overall, $48$ of $177$ missing components ($27\%$) were recovered through validation (Figure 6A).
-Incorporating the Pilot and Wave I, we re-fit the healthcare utilization model using the SMLE with the first $52$ validated and $948$ unvalidated patients’ data (Table 2).
+3.2.2 Wave II Chart Reviews
+In Wave II, $3867$ data points were validated ($718$ labs, $3149$ vitals), which was fewer than in the Pilot + Wave I.
+This difference was explained by fewer patient encounters, with $16$-$453$ validated data points per patient (median $=65.5$).
+The estimated error rates for non-missing components (TPR $=99\%$, FPR $=3\%$) and missing data recovery rate ($17\%$) in Wave II closely resembled the Pilot + Wave I (Figure 6B).
+Still, some auditors’ findings differed (Figure 5B).
+For example, more vitals matched Epic ($65\%$).
+Combined findings from all validation waves indicated that non-missing EHR data were of excellent quality (TPR $=99.5\%$, FPR $=2\%$), and auxiliary information could be located for $21\%$ of missing components (Figures 5C and 6C).
+On average, the number of non-missing components per patient increased from $6$ to $7$.
+Supplemental Tables S1–S2 contain detailed findings.
+3.3 Analysis Findings: Whole-Person Health and Healthcare Utilization
+Using ALI values from the EHR data, we fit first the naive healthcare utilization model.
+The expected odds of engaging in the healthcare system for an $18$-year-old patient with an ALI of 00 were $0.25$ (95% confidence interval [95% CI]: $0.18$, $0.36$).
+For every $0.1$-point (one-component) increase in ALI, these odds were expected to increase by $10\%$ (OR $=1.10$, 95% CI: $1.03$, $1.18$), adjusting for age.
+For every $10$ years older a patient was, their odds were expected to increase by $11\%$ (OR $=1.11$, 95% CI: $1.00,1.23$), adjusting for ALI.
+While these findings align with the literature, we wanted to ensure that missingness and potential errors in the EHR data did not lead to spurious conclusions.
+3.3.1 Preliminary Estimates Based on the Pilot + Wave I Validation
+Incorporating the Pilot + Wave I, we re-fit the healthcare utilization model with the first $52$ patients’ validated data and all $1000$ patients’ unvalidated data.
+These “preliminary SMLEs” were used to design the simulations in Section 2.6.1.
+Here, we compare them to the naive model.
+For an 18-year-old patient with an ALI of 00, the expected odds of engaging in the healthcare system decreased to $0.17$ (95% CI: $0.12$, $0.26$).
 The association between whole-person health and healthcare utilization appeared even stronger, with an estimated 21% increase in odds for each $0.1$-point increase in ALI (OR $=1.21$, 95% CI: $1.17,1.25$), adjusting for age.
-For an 18-year-old patient with an ALI of 00, the expected odds of engaging in the healthcare system decreased to $0.17$ (95% CI: $0.12$, $0.26$).
-The adjusted OR for age was relatively unchanged (OR $=1.10$, 95% CI: $1.01,1.21$).
-Thus, the preliminary SMLE model reaffirms that worse whole-person health is associated with higher expected odds of engaging in the healthcare system.
-Odds/Odds
-Coefficient Ratio Estimate 95% CI
-Naive Analysis (Unvalidated Data)
-Intercept $0.251$ $(0.176,0.357)$
-ALI $1.099$ $(1.028,1.175)$
-Age $1.108$ $(1.000,1.228)$
-Preliminary SMLE (Pilot + Wave I Validated Data)
-Intercept $0.173$ $(0.115,0.262)$
-ALI $1.207$ $(1.167,1.247)$
-Age $1.104$ $(1.007,1.211)$
-Final SMLE (All Waves Validated Data)
-Intercept $0.225$ $(0.156,0.327)$
-ALI $1.120$ $(1.088,1.154)$
-Age $1.105$ $(1.018,1.200)$
-Table 2: Estimates and 95% confidence intervals for coefficients from the naive analysis (logistic regression, no validated data), preliminary sieve maximum likelihood estimator (SMLE, validated data on $52$ patients), and final SMLE (validated data on $100$ patients) models.
-3.3.3 Simulations: Using Preliminary Audit Findings to Design Wave II
-Before validating the remaining patients, we designed bespoke simulation studies based on the preliminary audit data.
-These simulations allowed us to explore how potential designs compared under different scenarios for the true underlying data quality.
-In particular, the design offering the best statistical efficiency for the OR on ALI was desired.
-Focusing here on the takeaways and our decision-making, details like the data-generating mechanisms can be found in the Supplemental Materials.
-The five designs described in Section 2.3 were considered: SRS, CC, BCC, OPT, and RS.
-Under varied error rates, the targeted validation study designs offered empirically unbiased estimates and higher efficiency (i.e., smaller variance) than the uninformative SRS strategy.
-The RS design provided the best efficiency under extra low to moderate error settings; the preliminary validation data indicated that we had a perfect TPR $=1.00$ and negligible FPR $<0.01$.
-The SMLE $\hat{\beta}_{1}$ was empirically unbiased under full/high data recovery rates ($90-100\%$ missing data recovered through validation) and different validation study designs.
-While biased for moderate recovery or lower ($\leq 50\%$), the SMLE was still closer to the truth than the naive estimate was (i.e., based on the unvalidated EHR data).
-The RS design offered the highest efficiency, regardless of how much missing data was recovered.
-Thus, based on extensive simulations considering different error rates and data recovery scenarios, we concluded that the RS design was the most informative way to select the remaining $n_{2}=48$ patients for validation in Wave II.
-3.3.4 Wave II: Implementing the Targeted Validation Study Design
-For the RS design, each unvalidated patient’s residual was calculated based on the naive model estimates, following (4), as
-$r_{i}^{*}=Y_{i}-\frac{1}{1+\exp\left\{-\left(-1.383+0.945X_{i}^{*}+0.103Z_{i}\right)\right\}}.$
-Then, patients were ranked by $r_{i}^{*}$, and the $24$ patients with the smallest residuals and the $24$ patients with the largest residuals were selected for validation.
-Here, the RS design prioritized (i) patients with zero hospitalizations or ED visits but high predicted probabilities of having at least one and (ii) patients with at least one hospitalization or ED visit but high predicted probabilities of having zero.
-That is, patients whose observed $Y_{i}$ aligned well with the naive model’s prediction were not considered to be very informative and thus were deprioritized for validation.
-A total of $3867$ data points were validated in Wave II ($718$ lab and $3149$ vitals variables), which was fewer than in the Pilot and Wave I.
-Given that similar numbers of patients were validated, this difference was explained by fewer patient encounters, with between $16$ and $453$ validated data points per patient (median $=65.5$).
-Overall, the estimated error rates for non-missing components (TPR $=99\%$, FPR $=3\%$) and recovery rate for missing data ($17\%$) in Wave II closely resembled the preliminary ones (Figure 6B).
-Still, some auditors’ findings differed (Figure 5B).
-For example, more vitals were found to match Epic ($65\%$).
-Similar amounts of missing data were recovered using the audit roadmap, spread across all lab variables.
-Additional findings are in Supplemental Table S1.
-3.4 Post-Validation Model of Healthcare Utilization After Validation
-Combined findings from all waves indicated that the non-missing EHR data were of excellent quality (TPR $=99.5\%$, FPR $=2\%$), and auxiliary information could be located for $21\%$ of missing components (Figures 5C and 6C).
-Detailed findings are in Supplemental Table S2.
-We pooled validated data for the $n=100$ total validated patients with the unvalidated on all $N=1000$ patients from the original EHR data to estimate the SMLE once more (Table 2).
-All final model estimates fell between the naive analysis and preliminary SMLEs.
-For each $0.1$-point increase in ALI, a $12\%$ increase in odds of engaging in the healthcare system was expected (OR $=1.12$, 95% CI: $1.09,1.15$), adjusting for age.
-The expected odds of engaging in the healthcare system for an 18-year-old patient with an ALI of 00 was $0.23$ (95% CI: $0.16$, $0.33$).
-Again, older patients were expected to be more likely to engage in the healthcare system, controlling for ALI (OR $=1.11$, 95% CI: $1.02,1.20$).
+The adjusted OR $=1.10$ for age was relatively unchanged (95% CI: $1.01,1.21$).
+Thus, the preliminary SMLEs reaffirmed that worse whole-person health was associated with higher expected odds of engaging in the healthcare system, perhaps more strongly than the naive model indicated.
+3.3.2 Final Estimates Based on All Waves of Validation
+After completing all three waves, we used the $100$ patients’ validated data with all $1000$ patients’ unvalidated data to re-fit the model one last time.
+All final estimates fell between the naive model and preliminary SMLEs.
+The expected odds of engaging in the healthcare system for an 18-year-old patient with an ALI of 00 were $0.23$ (95% CI: $0.16$, $0.33$).
+For each $0.1$-point increase in ALI, a $12\%$ increase in odds was expected (OR $=1.12$, 95% CI: $1.09,1.15$), adjusting for age.
+Again, older patients were expected to be more likely to engage, controlling for ALI (OR $=1.11$, 95% CI: $1.02,1.20$).
 Additional technical details, including R code, are available in the Supplemental Materials.
-Ultimately, all three models agreed that worse whole-person health (i.e., higher ALI) was associated with a higher likelihood of being hospitalized or visiting the emergency department, adjusting for age.
-Our chart reviews uncovered very high quality of the non-missing EHR data and located auxiliary information for some missing data.
-Without validation, knowing whether we can trust the EHR data would not have been possible.
-Thus, while the validated and unvalidated models led to the same conclusions here, validating EHR data before using them in research is critical.
+3.4 Informing Implementation in the Learning Health System
+Ultimately, all models agreed that worse whole-person health (i.e., higher ALI) was associated with a higher likelihood of engaging in the healthcare system, adjusting for age.
+Chart reviews uncovered very high quality of non-missing EHR data and located auxiliary information for some missing data.
+Without validation, knowing whether we could trust the EHR data and whether the EHR-derived ALI was still a valid computable phenotype for whole-person health would not have been possible.
+Thus, while the models led to the same conclusions here, validating EHR data before using them in practice or research is critical.
+Our findings can now inform implementation strategies to incorporate the ALI into this learning health system.
+Having established ALI as a strong predictor of healthcare utilization, embedding it in the EHR is a critical next step.
+If ALI automatically prepopulated in patients’ electronic charts, clinicians would have easy access to it when making care decisions and researchers could extract it for further studies.
+This healthcare system previously implemented another score, the EFI, across the patient population aged $65+$.
+Building on that experience and encouraged by the ALI’s predictive ability, we plan to operationalize it for patients $18$ years and older.
+Of course, we want to embed the validated ALI, and only a small subset of patients can undergo chart reviews.
+Fortunately, in obtaining SMLEs for the healthcare utilization model, we robustly estimated the exposure error mechanism, $\Pr(X|X^{*},Z)$.
+When scaling up across the healthcare system, the estimated exposure error mechanism can be used to efficiently predict patients’ validated ALI from unvalidated EHR data (Figure 7).
+See the Supplemental Materials for details on calculating these predictions.
+Figure 7: Among patients with chart reviews (triangles), validated allostatic load index (ALI) could differ noticeably from the unvalidated version using the EHR data.
+For patients without chart reviews (squares), validated ALI could be robustly predicted using the estimated exposure error mechanism.
+The dashed line indicates the line of equality (i.e., where unvalidated and validated ALI are the same).
+The solid line is a loess smoother.
 4 Discussion
-Predicting healthcare utilization is a priority in the learning health system, with important implications at both the patient- and institution-levels.
-Establishing standardized, valid summary measures of whole-person health, like the ALI, associated with utilization and can scale up to the entire institution is an integral step.
+Predicting healthcare utilization is a priority in the learning health system, with important implications at both patient and institutional levels.
+Establishing standardized, valid measures of whole-person health (like ALI) that are associated with utilization and can scale up to the entire institution is an integral step.
 While EHR data contain variables to calculate these measures, missing data is a key challenge, and non-missing data are still prone to errors.
-Data validation or auditing, often through chart reviews, can gauge data quality, but this time-intensive process is usually only possible for a small subset of patients’ data.
+Data validation, often through chart reviews, can gauge data quality, but this time-intensive process is usually only possible for a small subset of patients.
 Also, existing validation protocols do not overcome missingness.
-Herein, we propose novel methods to maximize information gained from a partial validation study of $100$ patients from an EHR dataset of $1000$ through the statistical model, validation protocol, and targeted strategy to select patients.
+We propose novel methods to maximize information gained from a partial validation study of $100$ patients from an EHR dataset of $1000$ through the statistical model, validation protocol, and targeted strategy to select patients.
+We focused on estimating the association between ALI and healthcare utilization.
+Partial validation can also efficiently improve estimates of the ALI itself, in preparation for implementation in the EHR.
 The proposed analytical and design strategies were devised to ensure robust and efficient estimates.
@@
 Moreover, the targeted design strategies are tailored to the preliminary validation data and chosen statistical model, selecting patients most informative for the clinical question.
-Our enriched validation protocol builds on those in the literature (e.g., [41]) by incorporating a clinically-driven “roadmap” to locate auxiliary information about missing values.
-While our roadmap helped recover some missing ALI components, the clinical guidelines were only defined for missing values that were presumed to be a “yes.” If auditors were given additional guidance on what auxiliary information might mean that a missing value is a “no,” the missingness could be further reduced.
-Others have described missingness in the EHR as selection bias and sought to correct for it using approaches like inverse probability weighting [42].
+Our enriched validation protocol builds on the literature (e.g., [52]) by incorporating a clinically-driven “roadmap” to locate auxiliary information about missing values.
+The robust SMLEs can handle conditioning on error-prone $X^{*}$ and maybe one other covariate $Z$ in the error mechanism, due to the curse of dimensionality.
+We assumed that validated ALI depended only on its unvalidated version.
+However, if a more complicated error structure is expected, data pre-processing will be needed before using the SMLEs Tao2021.
+While our roadmap helped recover some missing ALI components, validated patients could still have incomplete data.
+Clinical guidelines were only defined for missing values presumed to be unhealthy.
+If auditors were given additional guidance on what auxiliary information might mean that a missing value is healthy, missingness could be further reduced.
+Others have described missingness in the EHR as selection bias and sought to correct for it using approaches like IPW [53].
 Applying such approaches to our post-validation data may reduce the impact of residual missingness.
-Further, EHR data are susceptible to “informative presence bias,” wherein the process through which patients engage in the healthcare system can bias associations [43, 44, 45].
-For example, patients seeking care more often may have more diagnoses.
+Alternatively, we could adopt other allostatic load calculations, like the index of cardiometabolic health (ICMH), which rely less on rare labs and may face fewer missing components [9].
+EHR data are susceptible to “informative presence bias,” wherein the process through which patients engage in the healthcare system can bias associations [54, 55, 56].
 Our decision to adapt the Seeman et al.
-(2001) [5] definition of the ALI to be the proportion of non-missing components could be considered a summary measure.
-In a recent review, Sisk et al.
-(2021) [46] found summary measures to be the most common approach to handling informative presence bias.
-We also used the enriched chart reviews to better understand the missingness.
+[7] definition of the ALI as the proportion of non-missing components could be considered a summary measure.
+A recent review found summary measures to be the most common approach to handling informative presence bias [57].
+Whole-person health scores like ALI are among many computable phenotypes currently applied to EHR data [16].
+Other phenotyping algorithms, including EHRshot [58], also seek to translate observable characteristics into measures of underlying patient states.
+Validation is a critical step in their development and typically involves chart reviews to obtain some “gold standard” labels [59, 60].
+When phenotyping with machine learning methods, this validation set generally needs to be an SRS.
+Meanwhile, our statistical modeling approach accommodates targeted sampling of the validation study.
+This flexibility creates opportunities to promote the efficiency of downstream analyses (e.g., associations between the phenotype and suspected risk factors).
 There are several interesting directions for future work.
-First, patients could have multiple labs and vitals during the study period.
-The averages of these values were used to calculate the ALI, and a longitudinal approach to modeling healthcare utilization wherein patients’ ALI can vary over time might provide more information.
+First, a longitudinal approach to modeling healthcare utilization, wherein patients’ ALI can vary over time, might provide more information.
 Second, auditors recorded their findings in a Microsoft Excel workbook.
-An electronic REDCap [47] interface could be built following published guidelines [48] for ease of use and accuracy.
-Third, one could use other machine learning techniques (e.g., random forest classification) and explore targeted validation strategies to improve predictive ability (e.g., accuracy) rather than statistical efficiency [49, 50].
+For ease of use and accuracy, an electronic REDCap [61] interface could be built following published guidelines [35].
+Third, targeted validation strategies could be used to improve predictive ability, rather than efficiency [62, 63].
 5 Conclusion
-EHR data provide an exciting opportunity to implement key measures, like whole-person health scores, in the learning health system.
-However, errors and missingness in these routinely collected data pose challenges in verifying and operationalizing such measures.
-Validation of even a subset of patients’ data is time- and cost-intensive but provides crucial insights into data quality.
+EHR data provide an exciting opportunity to implement key measures in the learning health system.
+However, errors and missingness in these data pose challenges in verifying and operationalizing such measures.
+Validation of even a subset of patients’ data is time- and cost-intensive, but provides crucial insights into data quality.
 Statistical and biomedical informatics methods can maximize information gained from such partial validation.

### v5

@@
 Incorporating validation data, statistical models indicated that worse whole-person health (higher ALI) was associated with higher odds of engaging in the healthcare system, adjusting for age.
-Conclusion: Targeted validation with an enriched protocol enabled us to ensure the quality and promote the completeness of the EHR.
-Findings from our validation study were incorporated into analyses as we operationalize the ALI as a scalable whole-person health measure predictive of healthcare utilization in the learning health system.
+Conclusion: Targeted validation with an enriched protocol can ensure the quality and promote the completeness of EHR data.
+Findings from our validation study were incorporated into analyses as we operationalize the ALI as a scalable whole-person health measure that predicts healthcare utilization in the learning health system.
 Keywords chart review  $\cdot$ computable phenotype  $\cdot$ data audits  $\cdot$ healthcare utilization  $\cdot$ measurement error  $\cdot$ missing data  $\cdot$ residual sampling  $\cdot$ two-phase designs
@@
 The advent of learning health systems [1, 2] signals a transition from a reactive (treatment-based) to a proactive (prevention-based) philosophy in healthcare.
-This transition could prevent future healthcare utilization through early intervention, improving health for the individual and providing cost savings to the institution.
+This transition could prevent future healthcare utilization through early intervention, improving individual health and providing cost savings to the institution.
 Having a standardized summary measure of health that is simple to compute and track could play a central role in disease prevention (saving lives) and reducing healthcare utilization (saving costs).
-Numerous data types exist within the electronic health record (EHR), including vital signs, demographics, diagnoses, procedures, and clinical notes.
-Easy access to a standardized and valid summary measure of whole-person health in the EHR would be advantageous for patients, clinicians, and researchers [3, 4].
-The electronic frailty index (EFI) serves this purpose but is only designed for those aged 65 and older [5].
-A whole-person health metric for adults aged $18+$ would enable system-wide longitudinal tracking of individual and population health, comparisons across geographies and hospitals, and integration of an important predictor in algorithms relevant to the enterprise (e.g., predicting readmissions).
-Individual physiological variables measure aspects of health; they do not provide a comprehensive assessment across various organ systems.
+Having easy access to a standardized summary measure of whole-person health in electronic health records (EHR) would be advantageous for patients, clinicians, and researchers [3, 4].
+The electronic frailty index (EFI) serves this purpose but only for those aged $65$ and older [5].
+A whole-person health metric for adults aged $18$ and older would enable system-wide longitudinal tracking of individual and population health, comparisons across geographies and hospitals, and integration of an important predictor in algorithms relevant to the enterprise (e.g., predicting readmissions).
+Individual physiological variables only measure aspects of health.
 The allostatic load index (ALI) is a composite measure of whole-person health, which reflects the multiple interrelated physiological regulatory systems that underlie healthy functioning [6, 7, 8, 9].
-It has been used to predict numerous health outcomes (e.g., [10, 11, 12, 13]) and measure racial disparities in health [14].
+It has been used to predict numerous health outcomes (e.g., [10, 11, 12, 13]) and measure racial health disparities [14].
 Because it is sensitive to short-term interventions, ALI is also an appealing “surrogate” outcome to test health interventions.
 Many ALI formulations exist.
-We focus on the calculation from [7], which computes ALI from ten components across three body systems (Table 1).
-The original numeric measurements (e.g., hemoglobin A1c [HbA1c]) are discretized at clinically-driven thresholds to define binary indicators of being at an unhealthy level (e.g., $=1$ if HbA1c $\geq 6.5$ and $=0$ if HbA1c $<6.5$).
-ALI is ultimately calculated as the sum of these ten components, with zero and ten representing the best and worst whole-person health, respectively (Supplemental Figure S1).
+We focus on the Seeman et al.
+(2001) [9] calculation, which computes ALI from ten components across three body systems (Table 1).
+The original numeric measurements (e.g., hemoglobin A1c [HbA1c]) are discretized at clinically-driven thresholds to define binary indicators of being at an unhealthy level (e.g., =1=1 if HbA1c $\geq 6.5$ and =0=0 otherwise).
+ALI is the sum of these ten components, with zero and ten representing the best and worst whole-person health, respectively (Supplemental Figure S1).
 Whole-person health is a latent concept influenced by genetics, environment, and other factors.
-Scores like the ALI incorporate components across body systems and yield “computable phenotypes” for it, concisely characterizing patients’ current health status with one-number summaries of observed data [15, 16].
+Scores like the ALI yield “computable phenotypes” for it, concisely characterizing patients’ current health status with one-number summaries of observed data [15, 16].
 The ALI’s ability to quantify whole-person health has been rigorously examined in controlled studies [6, 7, 9].
-However, before adopting the ALI in other settings, it must be revalidated, accounting for real-world data challenges.
+However, before adopting the ALI in other settings, it must be revalidated accounting for real-world data challenges.
 1.2 Data Quality Challenges in Electronic Health Records Data
 EHR data present a substantial opportunity to operationalize the ALI in the learning health system.
-The volume of routinely collected data in the EHR is steadily climbing, increasing the accessibility to clinically meaningful variables.
-Analyzing data that come at little to no additional cost for collection is particularly appealing to biomedical researchers, leading to a surge in the utilization of EHR data in many clinical domains (e.g., [17, 18, 19, 20]).
-However, despite clear advantages, EHR data are collected in the fast-paced world of clinical care, leading to missingness and errors.
+The volume of routinely collected EHR data is steadily climbing, increasing the accessibility to clinically meaningful variables.
+Analyzing data that come at little cost for collection is particularly appealing to researchers, leading to a surge in EHR data use in many clinical domains (e.g., [17, 18, 19, 20]).
+However, EHR data are collected in the fast-paced world of clinical care, leading to missingness and errors.
 Quality concerns remain a significant hurdle to the large-scale adoption of EHR in healthcare research [21, 22, 23, 24].
-In calculating the ALI, measurements may be (i) missing, making the component missing, or (ii) non-missing but mismeasured, potentially misclassifying the component.
+In calculating the ALI, measurements may be (i) missing, making the component missing, or (ii) non-missing but mismeasured, misclassifying the component.
 Consider a patient whose true HbA1c was $6.5\%$, so their corresponding component should equal one.
 We illustrate two scenarios for this patient’s data in Figure 1.
-In the first, HbA1c was measured but erroneously transposed to $5.6\%$.
-Based on the EHR, their HbA1c component would be misclassified as zero since $5.6\%$ falls into the healthy range.
-In the second, HbA1c was not measured, so the value is missing, and we cannot define a component.
+First, suppose HbA1c was measured but erroneously transposed to $5.6\%$; their component would be misclassified as zero, since $5.6\%$ falls into the healthy range.
+Second, suppose HbA1c was not measured, so the value is missing.
+Then, we cannot define their component.
 Counting missing components as zeros would also induce misclassification by treating them as “healthy." We accommodate missing data by instead defining ALI as the proportion of non-missing components a patient experienced.
-Figure 1: Diagram of potential errors and missingness in a patient’s hemoglobin A1c (HbA1c) measurement in the electronic health record (EHR) before and after validation (chart review).
+Figure 1: Diagram of potential errors and missingness in a patient’s hemoglobin A1c (HbA1c) measurement in the electronic health records (EHR) before and after validation (chart review).
 In the top scenario, the patient had HbA1c measured but with error, and the correct value was recovered through validation.
 In the bottom scenario, the patient did not have HbA1c measured, but auxiliary information was located through validation to replace the missing component.
-Ignoring errors and missingness in the EHR can hinder analyses, as biased estimates may lead to invalid inference and incorrect conclusions [25, 26].
-Errors in EHR data can also reduce the accuracy of prediction models [27].
-There is a particular concern for missingness and errors in derived variables, like ALI, that combine data from many different variables [28, 29].
-Analyzing the ALI computed from the EHR data “as is” may misrepresent a patient’s actual health and lead to false conclusions about its association with healthcare utilization.
+Ignoring errors and missingness can hinder analyses, biasing inference [25, 26] and reducing predictive accuracy [27].
+There is particular concern for derived variables, like ALI, that combine multiple variables [28, 29].
+Analyzing the ALI computed from EHR data “as is” may misrepresent a patient’s actual health and lead to false conclusions about its association with healthcare utilization.
 Applying popular missing data approaches like multiple imputation (MI) may offer some improvement at baseline [30].
 However, they can also hinder results if any imputation models are misspecified [31].
-Since components are likely missing not at random in EHR data (e.g., clinicians usually order labs when they suspect values are abnormal) [32], correctly specifying imputation models could be especially challenging.
-Plus, MI would not correct for errors in non-missing ALI components.
-For the ALI to be useful in patient care and research, it must be reliably and accurately measured.
-The original calculations, like [7], were done in clinical trials, where all patients must have all measurements, and data validation protocols are likely in place to ensure quality.
+Since components are likely missing not at random in EHR data (e.g., clinicians usually order labs when they suspect abnormal values) [32], correctly specifying imputation models could be especially challenging.
+Plus, MI would not correct for errors.
+For the ALI to be useful, it must be reliably and accurately measured.
+The original calculations, like Seeman et al.
+(2001) [9], were done in clinical trials, where all patients must have all measurements.
+Further, data validation protocols were likely in place to ensure quality.
 Similarly, validation via chart review, comparing extracted EHR data to source documents (like electronic charts), provides data quality insights.
-Complete data validation (for all patients and variables) would ensure the integrity of EHR data but is overly expensive and unattainable in practice.
-Fortunately, partially validating EHR data can help identify issues and pave the way for corrections.
+Complete data validation (for all patients and variables) would ensure the integrity of EHR data but is expensive and unattainable.
+Fortunately, partially validating EHR data can identify issues and pave the way for corrections.
 It provides a cost-effective alternative to complete data validation and is a promising approach to gauge the quality of observational data sources [33, 34, 35].
-These partial validation studies are a special case of two-phase designs [36, 37].
-Phase I includes EHR data for all $N$ patients, which can be used to select a subsample of $n$ patients ($n<N$) for validation (chart reviews) in Phase II.
+Partial validation studies are two-phase designs [36, 37].
+Phase I includes EHR data for $N$ patients, which can be used to select a subsample of $n$ patients ($n<N$) for validation (chart reviews) in Phase II.
 Validated data are usually expected to be more correct (i.e., less error-prone) but not more complete (i.e., less missingness).
@@
 System Component Threshold Auxiliary Information
-Cardiovascular Systolic Blood Pressure $>140$ Hypertension
-Diastolic Blood Pressure $>90$ Hypertension
-Metabolic Body Mass Index $>30$ Obesity; Morbid Obesity;
+Cardiovascular Systolic Blood Pressure >140>140 Hypertension
+Diastolic Blood Pressure >90>90 Hypertension
+Metabolic Body Mass Index >30>30 Obesity; Morbid Obesity;
 Grade I, II or III Obesity
@@
 Serum Albumin $\geq 3.5$ (None Given)
-Creatinine Clearance $<110$ (Males) Renal Failure;
-$<100$ (Females) Insufficiency; Acute
+Creatinine Clearance <110<110 (Males) Renal Failure;
+<100<100 (Females) Insufficiency; Acute
 Kidney Injury;
 Chronic Renal Failure
-Homocysteine $>50$ Hyperhomocysteinemia;
+Homocysteine >50>50 Hyperhomocysteinemia;
 Vitamin deficiency
-Table 1: Ten components of the allostatic load index were defined by discretizing measurements across three body systems at clinically-driven thresholds.
-If the component was missing from the extracted EHR data, auditors searched for auxiliary information in the patient’s medical chart in Epic (the institution’s electronic charting software).
-If the auxiliary information was present, the patient was treated as having an “unhealthy” measurement and the component included in their validated ALI.
+Table 1: Ten components of the allostatic load index (ALI) were defined by discretizing measurements across three body systems at clinically-driven thresholds.
+If the component was missing from the extracted electronic health records (EHR) data, auditors searched for auxiliary information in the patient’s medical chart in Epic (the institution’s electronic charting software).
+If the auxiliary information was present, the patient was treated as having an “unhealthy” measurement and the component was included in their validated ALI.
 1.3 Application to the Learning Health System
 We investigate how to overcome data challenges to measure whole-person health across one of the largest learning health systems in the United States.
-Having recently integrated the EFI into the EHR to measure whole-person health for patients $65$ and older [38], the health system is interested in operationalizing the ALI for the broader patient population.
+Having recently integrated the EFI into the EHR to measure whole-person health for patients aged $65$ and older [38], the health system is interested in operationalizing the ALI for the broader patient population.
 We obtained a representative sample of $N=1000$ patients from the EHR at Atrium Health Wake Forest Baptist Hospital in Winston-Salem, North Carolina.
-Our sample was drawn from patients (i) $18$-$65$ years old who (ii) initiated their first primary care outpatient encounter in the health system between March 11, 2018, and March 10, 2020 (to avoid data challenges and biases associated with the COVID-19 pandemic).
-Outpatient encounters were chosen to represent the typical patient voluntarily engaging in care.
-Our data contained demographics, vitals, lab values, discharge diagnoses, and problem lists dating $3$ years before the first encounter.
-Key variables included measurements for the ten ALI components (Table 1), age at first encounter (hereafter, “age”), and an indicator of having at least one hospitalization or ED visit during the $2$-year study period.
+Our sample was drawn from patients (i) 18186565 years old who (ii) initiated their first primary care outpatient encounter in the health system between March 11, 2018 and March 10, 2020.
+Outpatient encounters were chosen to represent patients voluntarily engaging in care.
+Our data contained demographics, vitals, and lab values dating back $3$ years before the first encounter.
+Key variables included measurements for the ten ALI components (Table 1), age at first encounter (hereafter, “age”), and an indicator of having at least one hospitalization or emergency department (ED) visit during the $2$-year study period.
 Since most patients had multiple encounters, ALI components were based on mean measurements.
@@
 Using a representative sample of $1000$ patients from the EHR at an extensive academic learning health system ($100$ of whom could be validated), we propose novel methods to design, conduct, and analyze statistically efficient and robust studies of whole-person health (ALI) and healthcare utilization.
-Using advanced statistics and bioinformatics methods, we seize three key opportunities to maximize our investment into partial validation: robust and efficient statistical modeling, enriched chart review protocol, and targeted sampling design.
+Using statistics and biomedical informatics methods, we seize three key opportunities to maximize our investment into partial validation: robust and efficient statistical modeling, enriched chart review protocol, and targeted sampling design.
 Statement of Significance
-  • Problem Standardized and valid summary measures of current health, i.e., whole-person health, are needed in electronic health records (EHR) to make patient- and institution-level decisions.
+  • Problem Standardized summary measures of whole-person health are needed in the EHR to make patient- and institution-level decisions.
 However, routinely collected data are prone to missingness and errors.
-  • What Is Already Known Validation, wherein the extracted EHR data are compared to the original clinical source documents (like the electronic chart), provides insights into data quality, but validating all patients is not feasible, and existing protocols can only correct for errors, not missingness.
-  • What This Paper Adds Building on the statistics and bioinformatics literature, we 1) devise a robust and efficient model for partially validated, error-prone EHR data; 2) propose a targeted validation study design to identify the most informative patients for our model; and 3) develop an enriched validation protocol that incorporates clinical expertise to improve the quality and completeness of EHR data.
-  • Who Would Benefit From the Knowledge in This Paper Biomedical researchers interested in analyzing routinely collected data, like those in the EHR, while correcting their data quality issues.
+  • What Is Already Known Validation, wherein extracted EHR data are compared to clinical source documents (like electronic charts), provides insights into data quality, but validating all patients is not feasible, and existing protocols only correct for errors, not missingness.
+  • What This Paper Adds We devise a robust and efficient model for partially validated, error-prone EHR data; propose a targeted validation study design to identify the most informative patients for our model; and develop an enriched validation protocol that incorporates clinical expertise to improve the quality and completeness of EHR data.
 2 Methods
@@
 This study was approved by the Institutional Review Board at Wake Forest University School of Medicine (WFUSM).
-We model healthcare utilization ($Y\in\{0,1\}$), defined as a binary indicator of having at least one hospitalization or emergency department (ED) visit in the $2$-year study period.
+We model healthcare utilization ($Y\in\{0,1\}$), defined as a binary indicator of having at least one hospitalization or ED visit in the $2$-year study period.
 Whole-person health, measured by ALI ($X\in[0,1]$), is the exposure of interest.
@@
 Further, we control for age ($Z\in\mathcal{R}$).
-We assume a logistic regression model for the log-odds that $Y=1$ (patient engaged in healthcare system) given ($X$, $Z$).
-A sample of $N$ patients was extracted from the EHR to estimate these log odds ratios (logORs), denoted ${\boldsymbol{\beta}}$.
-Age $Z$ is fully observed and expected to be error-free.
-However, due to missingness and errors in ALI components, only an error-prone measure for $X$, denoted by $X^{*}$ ($X^{*}\in[0,1]$), is available from the EHR data.
+We assume a logistic regression model for the log-odds that $Y=1$ given ($X$, $Z$).
+A sample of $N$ patients was extracted from the EHR to estimate these log odds ratios (logORs), denoted by $\boldsymbol{\beta}$}.
+Age $Z$ is fully-observed and expected to be error-free.
+However, only an error-prone measure for $X$, denoted by $X^{*}$ ($X^{*}\in[0,1]$), is available from the EHR data.
 We define error-prone components $S_{j}^{*}$ ($j\in\{1,\dots,10\}$) and introduce missingness indicators $M_{j}^{*}$ for them (i.e., $M_{j}^{*}=1$ if $S_{j}^{*}$ is missing).
 We compute $X^{*}=\sum_{j=1}^{10}(1-M_{j}^{*})S_{j}^{*}/\sum_{j=1}^{10}(1-M_{j}^{*})$, i.e., as the proportion of non-missing components experienced.
-Calculating ALI as a proportion – rather than sum – was critical.
-Summing over non-missing components treats missing values as zeros (i.e., healthy measurements); instead, the proportion ignores them.
+Calculating ALI as a proportion was critical.
+The sum of non-missing components treats missing values as zeros, whereas the proportion ignores them.
 2.2 Three Key Steps to Partial Validation Studies: Analysis, Design, and Protocol
-To obtain consistent estimates for the healthcare utilization model, we planned a partial validation study.
-We first extracted EHR data ($Y$, $X^{*}$, $Z$) for a sample of $N=1000$ patients.
-Then, a subset of $n=100$ patients was selected for validation via chart review such that $X$ was also measured.
+To consistently estimate the healthcare utilization model, we planned a partial validation study.
+We extracted EHR data ($Y$, $X^{*}$, $Z$) for a sample of $N=1000$ patients.
+A subset of $n=100$ patients was selected for validation (chart review) such that $X$ was also measured.
 Partial validation studies break down into three key steps.
-Design: Select which patients should be validated.
-Protocol: Develop standardized tools and practices for those patients’ chart reviews.
-Analysis: Estimate statistical models or quantities of interest from validated and unvalidated data together.
+Design: We select which patients should be validated.
+Protocol: We develop standardized tools and practices for those patients’ chart reviews.
+Analysis: We estimate statistical models from validated and unvalidated data together.
 These steps are intertwined, with design choices depending on the planned analysis, and each step provides opportunities to gain information.
-For added flexibility, a multi-wave validation framework, conducting chart reviews in multiple sprints, allows us to revisit key steps and make modifications as information accumulates (Figure 2).
-In our application (Section 3), we built the statistical model and developed the baseline sampling strategy and tools before beginning validation.
-The proposed methods are detailed in order of development: analysis (Section 2.3), design (Section 2.4), and protocol (Section 2.5).
-Across three validation waves, we updated our sampling strategy and tools (revisiting the design and protocol steps, respectively), before fitting the final model (Section 2.6).
-Figure 2: After extracting the electronic health records (EHR) data, partial validation studies break down into three key steps: analysis, design, and protocol.
-These steps are intertwined, with design choices depending on the planned analysis, and each step provides opportunities to gain information.
-For added flexibility, a multi-wave validation framework allows us to revisit these key steps and make modifications as information accumulates.
+For added flexibility, a multi-wave validation framework (i.e., conducting chart reviews in multiple sprints) allows us to revisit key steps and make modifications as information accumulates (Figure 2).
+We built the statistical model and developed the baseline sampling strategy and tools before beginning validation.
+The proposed methods are detailed in order of development: analysis, design, and protocol.
+Across three validation waves, we updated our sampling strategy and tools (revisiting the design and protocol steps, respectively) before fitting the final model.
+Figure 2: After extracting the electronic health records (EHR) data, we prepared for validation by defining our analysis plan, picking an initial sampling strategy, and developing baseline validation tools.
+We first conducted the Pilot Wave chart reviews.
+This information was then used to modify the validation protocol and establish inter-auditor agreement.
+Next, the Wave I chart reviews followed the same sampling strategy as the Pilot but used the updated validation tools.
+With the EHR data ($N=1000$) and chart review data from the Pilot and Wave I ($n=52$), we fit a preliminary version of the healthcare utilization model.
+Estimates from this model guided our selection of the targeted sampling strategy for Wave II via bespoke simulation studies.
+Following completion of all validation waves, we combined the EHR data ($N=1000$) with all the chart review data ($n=100$) to fit the final healthcare utilization model.
 2.3 Analysis: Robust and Efficient Statistical Model for Partially Validated Data
-Let $V_{i}$ distinguish validated and unvalidated patients, where $V_{i}=1$ if the $i$th patient is validated, $i\in\{1,\dots,N\}$, and $V_{i}=0$ otherwise.
+Let $V_{i}$ distinguish validated and unvalidated patients, where $V_{i}=1$ if patient $i$ is validated ($i\in\{1,\dots,N\}$) and $V_{i}=0$ otherwise.
 Incorporating EHR and chart review data, along with the validation indicator, the joint distribution of a complete observation is assumed to be
@@
 Two assumptions are made in (1).
-First, selection for validation depends only on fully observed variables ($Y$, $X^{*}$, $Z$), so the validated exposure $X$ is missing at random (MAR) for unvalidated patients [39].
-Second, $X^{*}$ from the EHR is a surrogate for validated $X$ from the chart reviews, such that $Y$ and $X^{*}$ are conditionally independent given $X$.
+First, validation sampling depends only on fully-observed variables ($Y$, $X^{*}$, $Z$), so $X$ is missing at random (MAR) for unvalidated patients [39].
+Second, $X^{*}$ is a surrogate for $X$, such that $Y$ and $X^{*}$ are conditionally independent given $X$.
 Observations for all $N=1000$ patients are assumed to be independent and identically distributed following (1).
-Thus, the observed-data log-likelihood function (hereafter, log-likelihood) for the logORs ${\boldsymbol{\beta}}$ is defined from (1), after two simplifications.
+Thus, the observed-data log-likelihood function (hereafter, log-likelihood) for the logORs $\boldsymbol{\beta}$} is defined from (1) after two simplifications.
 First, we can omit $\Pr(V|Y,X^{*},Z)$ from the log-likelihood because $X$ is MAR.
-Second, we can ignore $\Pr(X^{*},Z)$ because it drops out as the log-likelihood is maximized with respect to ${\boldsymbol{\beta}}$ since ($X^{*}$, $Z$) are fully observed.
+Second, we can ignore $\Pr(X^{*},Z)$ because it drops out as the log-likelihood is maximized with respect to $\boldsymbol{\beta}$}, since ($X^{*}$, $Z$) are fully-observed.
 Now, the log-likelihood for $\boldsymbol{\beta}$} is proportional to
@@
 $\sum_{x=0}^{1}{\Pr}_{{\boldsymbol{\beta}}}(Y_{i}\mid x,Z_{i})\Pr(x\mid X_{i}^{*},Z_{i})$ $\propto\sum_{x=0}^{1}{\Pr}_{{\boldsymbol{\beta}}}(Y_{i},x,Z_{i},X_{i}^{*})={\Pr}_{{\boldsymbol{\beta}}}(Y_{i},Z_{i},X_{i}^{*}),$ (3)
-remembering that $X$ takes on values $x\in\{0,0.1,\dots,0.9,1\}$ here.
-A model for $\Pr(X\mid X^{*},Z)$ is needed to estimate ${\boldsymbol{\beta}}$ from (2) since $X$ is only observed for validated patients.
-Since $X\in[0,1]$, there are not many flexible, parametric choices to model its distribution given ($X^{*},Z$).
+where $X$ takes on values $x\in\{0,0.1,\dots,0.9,1\}$ here.
+A model for $\Pr(X\mid X^{*},Z)$ is needed to estimate $\boldsymbol{\beta}$} from (2), since $X$ is only observed for validated patients.
+Since $X\in[0,1]$, there are not many flexible, parametric choices to model this conditional distribution.
 A nonparametric estimator is preferable, allowing $X$ to take on any values and arbitrarily depend on ($X^{*}$, $Z$).
-[40] proposed a semiparametric approach to logistic regression with outcome and exposure errors, which nonparametrically (i.e., without assuming a particular distribution or structure) estimates the exposure error mechanism as described.
-The current setting is a special case with only exposure error, and a brief outline is provided in the Supplemental Materials.
+Lotspeich et al.
+(2022) [40] proposed semiparametric sieve maximum likelihood estimators (SMLEs) for logistic regression with outcome and exposure errors.
+They nonparametrically handle the exposure error mechanism as described.
+A brief outline of how to adapt the SMLEs for exposure error only is provided in the Supplemental Materials.
 The logiSieve package contains R functions to fit this model [41].
-For full technical details on the semiparametric sieve maximum likelihood estimators (SMLEs), see [40].
 2.4 Design: Targeted Sampling to Select Informative Patients for Validation
@@
 However, for full-likelihood approaches like SMLEs, any information available for all patients can be used to design the validation study.
-Designing based on fully observed information ensures that the MAR assumption holds for the downstream analysis while allowing for many targeted strategies.
-Figure 3: Diagram of partial validation studies with $N$ patients extracted from the electronic health records (EHR) and $n$ patients chosen for validation via chart review.
+Designing based on fully-observed information ensures that the MAR assumption holds for the downstream analysis while allowing for many targeted strategies.
+Figure 3: Illustrative diagram of partial validation studies with $N=48$ patients extracted from the electronic health records (EHR) and $n=16$ patients chosen for validation via chart review.
 The least informative strategy to select the validation study is A) simple random sampling.
-More informative targeted strategies include B) case-control sampling, C) balanced case-control sampling using discretized ALI, D) optimal stratified sampling using discretized ALI, and E) residual sampling based on the naive model.
+More informative targeted strategies include B) case-control sampling, C) balanced case-control sampling using discretized ALI, D) optimal stratified sampling using discretized ALI, and E) residual sampling based on the naive healthcare utilization model.
 Within strata, the darker shaded areas indicate which patients would be selected.
-When designing a validation study to promote statistical efficiency of the parameter estimates (i.e., to minimize the variance in the estimated logOR on ALI), we start at the end.
+When designing a validation study to promote statistical efficiency of a parameter estimate (e.g., to minimize the variance of the estimated logOR $\hat{\beta}_{1}$ on ALI), we start at the end.
 Beginning with the planned logistic regression model of $Y$ given ($X$, $Z$), we work backward to determine which patients will be most informative.
-For this reason, the “best” designs vary across different estimators, with model-based ones, like SMLEs, offering the most design flexibility.
-Estimators that rely on the design to handle missingness, like inverse probability weighting (IPW), require nonzero sampling probability for all patients; model-based estimators do not.
+For this reason, the “best” designs vary across different estimators, with model-based ones (like the SMLEs) offering the most flexibility.
+Estimators that rely on the design to handle missingness (like inverse probability weighting [IPW]) require a nonzero sampling probability for all patients; model-based estimators do not.
 We considered stratified and rank-based designs.
-Multiple validation waves were conducted to allow the protocol and study design to adapt as information accumulated, which is becoming common in this space [42, 34, 43].
-Ultimately, simulations based on preliminary validation data guided our selection of the targeted design used in the application (see Section 2.6).
+Multiple validation waves were conducted to allow the protocol and study design to adapt as information accumulated, which is becoming common practice in this space [42, 43, 35].
+Ultimately, simulations based on preliminary validation data (Section 2.6) guided our selection of the targeted design used in the application.
 2.4.1 Stratified Designs
-Given our binary outcome, the case-control (CC) design is a natural option, wherein $50$ patients each are chosen from subgroups with $Y=0$ and $Y=1$ (Figure 3B).
-The balanced case-control (BCC) design further stratifies on covariates, subsequently choosing equal (i.e., balanced) numbers from each stratum (Figure 3C).
+Given our binary outcome, the case-control (CC) design is a natural option, wherein $50$ patients each are chosen with $Y=0$ and $Y=1$ (Figure 3B).
+The balanced case-control (BCC) design further stratifies on covariates and chooses equal (balanced) numbers from each stratum (Figure 3C).
 Continuous covariates must be discretized to form strata.
-For instance, we could use $X_{D}^{*}$, a discretized version of $X^{*}$, to define strata of patients with error-prone ALI above and below the average.
-Then, validate $25$ patients from each of the four ($Y$, $X_{D}^{*}$) strata.
+For instance, we could let $X_{D}^{*}$ define strata of patients with error-prone ALI $X^{*}$ above and below the average and then validate $25$ patients from each of the four ($Y$, $X_{D}^{*}$) strata.
 Importantly, we can still fit the model using the continuous values $X^{*}$.
 The CC and BCC designs are intuitive and offer efficiency gains over SRS.
-Still, slightly more complicated designs are guaranteed to be theoretically “optimal;” they choose the number of patients from each stratum to achieve the lowest possible variance for the logOR on ALI [44, 45, 46, 47, 37, 48].
-[43] proposed the optimal (OPT) design for model-based analysis of data with binary outcome and exposure misclassification (Figure 3D), which could apply here.
+Still, slightly more complicated designs are guaranteed to be theoretically “optimal;” they choose the number of patients per stratum to achieve the lowest possible variance for $\hat{\beta}_{1}$ [44, 45, 46, 47, 48, 36].
+Lotspeich et al.
+(2024) [42] proposed the optimal (OPT) design for model-based analyses with binary outcome and exposure misclassification (Figure 3D), which could apply here.
 However, we would achieve the lowest possible variance for $X_{D}^{*}$ rather than $X^{*}$.
-Increasing the number of unique values of $X_{D}^{*}$ (i.e., more strata) can help, but information will still be lost.
-Moreover, the OPT design depends on the true logORs, which are unknown; reliable placeholders for them are required to implement the design (often collected in a multi-wave framework).
+Increasing the number of unique values of $X_{D}^{*}$ (i.e., creating more strata) can help, but information will still be lost.
+Moreover, the OPT design depends on the true logORs $\boldsymbol{\beta}$}, which are unknown.
+Reliable placeholders for $\boldsymbol{\beta}$} (often collected in a multi-wave framework) are required to implement the design.
 2.4.2 Rank-Based Designs
-Designs that rank patients rather than stratifying them offer improved granularity and sizable efficiency gains.
-For example, extreme tail sampling (ETS) designs rank patients by some value of interest and target those with the smallest or largest (i.e., most “extreme”) values.
-To identify patients most informative not just about ALI but its relationship to healthcare utilization, residual sampling (RS) is preferred [49].
-To implement RS with measurement error, we begin by fitting the “naive” analysis model, i.e., using standard logistic regression procedures with $Y$, $X^{*}$ (in place of $X$), and $Z$ to obtain estimates for the logORs $\hat{\boldsymbol{\beta}}^{*}$ from the full sample of $1000$ patients.
-Then, the naive model residuals are computed for each patient $i$ as
+Designs ranking patients offer improved granularity and efficiency gains.
+For example, extreme tail sampling (ETS) designs order patients by some value of interest and target those with the smallest or largest (most “extreme”) values.
+To identify patients who are most informative to the healthcare utilization model, residual sampling (RS) is preferred [49].
+We begin by fitting the “naive” analysis model, i.e., using standard logistic regression procedures with $Y$, $X^{*}$ (instead of $X$), and $Z$ to estimate the logORs $\hat{\boldsymbol{\beta}}^{*}$ from the full sample.
+Then, we compute residuals for all patients as
 $r_{i}^{*}$ $=Y_{i}-{\Pr}_{\hat{\boldsymbol{\beta}}^{*}}(Y_{i}=1|X^{*}_{i},Z_{i})$
 $=Y_{i}-\left[1+\exp\left\{-\left(\hat{\beta}_{0}^{*}+\hat{\beta}_{1}^{*}X_{i}^{*}+\hat{\beta}_{2}^{*}Z_{i}\right)\right\}\right]^{-1},$ (4)
-and used to rank patients.
-The largest residuals $r_{i}^{*}$ (in magnitude) correspond to patients whose observed outcome $Y_{i}$ had a low predicted probability (Figure 3E).
+and sample the $100$ patients with the most extreme residuals (Figure 3E).
+The largest residuals (in magnitude) correspond to patients whose observed $Y_{i}$ had a low predicted probability.
 The auditDesignR package contains R functions to sample from all of these designs [50].
 2.5 Protocol: Enriched Validation Protocol to Reduce Errors and Missingness
-Chart reviews were conducted over $6$ months by four clinical research assistants and technicians at WFUSM (hereafter, “auditors”), chosen for their familiarity with Epic (the institution’s electronic charting software) and demonstrated research experience.
+Chart reviews were conducted over $6$ months by four clinical research technicians at WFUSM (hereafter, “auditors”), who were chosen for their familiarity with Epic (the institution’s electronic charting software) and research experience.
 2.5.1 Identifying Sources of Auxiliary Information
-Before the chart reviews, our clinical collaborators developed the “audit roadmap” (Table 1).
-While it is unlikely that we would identify the original missing measurements, we harnessed clinical expertise to identify auxiliary information about missing discretized components.
+Before the chart reviews, we developed the “audit roadmap” (Table 1), harnessing clinical expertise to identify auxiliary information about missing ALI components.
 The roadmap suggested where auditors could look in Epic to locate supplemental information about missing values.
-It employs “anchors” to try and recover missing components using clinically-driven placeholders [15].
-For example, if a patient were missing HbA1c, auditors would search their chart for a diagnosis of diabetes or impaired glycemic control.
+It employs “anchors” to try to recover missing components using clinically-driven placeholders [15].
+For example, if a patient was missing HbA1c, auditors would search their chart for a diabetes or impaired glycemic control diagnosis.
 If either was present, the patient’s diagnosis would suggest that their missing HbA1c value would have exceeded the $6.5\%$ threshold.
-Thus, their validated ALI $X$ would include one point for HbA1c.
-If neither diagnosis was present, $X$ would continue to exclude this still-missing component.
+Thus, their validated $X$ would include one point for HbA1c.
+If neither diagnosis was present, $X$ would exclude this still-missing component.
 Many components, like HbA1c, had multiple sources of auxiliary information, while others, like triglycerides, had just one.
-Only serum albumin, a blood test to measure liver and kidney function, did not have a supplemental diagnosis.
+Only serum albumin, a blood test for liver and kidney function, did not have any.
 2.5.2 Validated Data Collection
-In preparation for validation, the analytical team transformed the extracted EHR data from the original wide format (one row per patient encounter and columns for different variables) to a long format (one row per patient encounter per variable).
-This transformation streamlined the data structure for auditors: one value to review per row (Supplemental Figure S2).
-Before sending the data to the auditors, the roadmap was incorporated, and an empty column was added for the “reviewed value” (i.e., the one in Epic).
+In preparation, the extracted EHR data were transformed from a wide (one row per patient encounter) to a long format (one row per patient encounter per variable).
+This transformation streamlined the data structure for auditors, with one value per row to review (Supplemental Figure S2).
+Before sending the data to the auditors, we incorporated the roadmap and added an empty column for the “reviewed value” (i.e., the one in Epic).
 Depending on the scenario, auditors would enter different information for the reviewed value, and each data point was assigned one of five auditor findings (Figure 4).
@@
 2.6 Multi-Wave Strategy: Sampling Adaptively and Improving Tools
-Choosing the “best” design often relies on information that is unknown at baseline, like EHR error rates or the magnitude of the association between ALI and healthcare utilization.
-Since we did not have any prior information about these unknowns, we started by sampling conservatively to collect preliminary information for a more extreme, targeted design later on.
-Multi-wave validation schemes like this one are appealing because they balance learning about our data with strategically designing for them [42, 34].
-Even if prior information is available (e.g., from previous chart reviews), multi-wave validation allows us to relearn how data have changed before dedicating resources to validate them [43].
-Further, we can seize opportunities to incorporate auditors’ feedback and improve the validation protocol.
+Choosing the “best” design often relies on information that is unknown at baseline, like the EHR error rates or the logORs for the healthcare utilization model.
+Since we did not have any prior information about these unknowns, we started by sampling conservatively to collect information for a targeted design later.
+This multi-wave validation scheme balances learning about our data with strategically designing for them.
+Even if prior information is available (e.g., from previous chart reviews), multi-wave validation allows us to learn how our data have changed before dedicating resources to validate them again [42].
+Further, we can incorporate auditors’ feedback between waves to improve the validation protocol.
 We validated $100$ patients’ data across three waves: the Pilot, Wave I, and Wave II.
-The small Pilot wave tested the validation protocol and established inter-auditor agreement on the first four patients.
-Then, Waves I and II put this protocol into action, as auditors validated the remaining patients.
+The Pilot wave tested the validation protocol and established inter-auditor agreement on the first four patients.
+Then, Waves I and II put this protocol into action as auditors validated the remaining patients.
 Waves I and II each selected $48$ more patients.
-(Designating equal-sized waves balances design flexibility and stability [42].) In Wave II, information from the Pilot + Wave I was leveraged to design a more targeted strategy.
+(Designating equal-sized waves balances design flexibility and stability [43].) In Wave II, information from the Pilot and Wave I was leveraged to design a more targeted strategy.
 Patients could be chosen for validation at most once.
-2.6.1 Adapting the Validation Study Design as Information Accumulates
-Between the Pilot and Wave I, the first $52$ patients were selected for chart review via the BCC design.
+2.6.1 Modifying the Validation Study Design as Information Accumulates
+Between the Pilot and Wave I, $52$ patients were selected for chart review via the BCC design.
 Specifically, patients were split into four strata based on their outcome and error-prone ALI (discretized at the median of $0.33$).
 Then, equal numbers of patients were randomly chosen from the four strata.
-In the Pilot, one patient was chosen from each stratum; another $12$ were chosen per-stratum in Wave I.
-Before validating the remaining $48$ patients, we designed bespoke simulation studies based on the preliminary validation data.
-They explored how designs compared under different scenarios for true underlying data quality.
-The design offering the best efficiency (i.e., lowest variability) in estimating the logOR on ALI with the SMLE was desired.
-Under varied error rates, the targeted validation study designs offered empirically unbiased estimates and higher efficiency than SRS.
-The RS design provided the best efficiency under extra low to moderate error settings, and the preliminary validation data (Pilot + Wave I) indicated a perfect true positive rate (TPR) $=1.00$ and negligible false positive rate (FPR) $<0.01$.
-Under full/high data recovery rates ($90$-$100\%$ missing data recovered through validation) and different designs, the SMLE was empirically unbiased.
-While biased for moderate recovery or lower ($\leq 50\%$), the SMLE was still closer to the truth than the naive estimate was (i.e., based on unvalidated EHR data).
-Regardless of how much missing data was recovered, the RS design offered the highest efficiency.
-Details like the data-generating mechanisms can be found in the Supplemental Materials.
+In the Pilot, one patient was chosen from each stratum; in Wave I, another $12$ were.
+Before beginning Wave II, we designed simulation studies based on the preliminary validation data to explore how designs compared under different scenarios for true underlying data quality.
+The design offering the best efficiency (i.e., lowest variability) for the SMLE $\hat{\beta}_{1}$ of the logOR on ALI was desired.
+Details are in the Supplemental Materials.
+Under varied error rates, the targeted designs offered empirically unbiased estimates and higher efficiency than SRS.
+The RS design provided the best efficiency under extra-low to moderate error settings, and the preliminary validation data (Pilot and Wave I) indicated a true positive rate (TPR) =1.00=1.00 and a false positive rate (FPR) <0.01<0.01.
+[…]
 We concluded that the RS design was the most informative way to select the remaining $48$ patients for validation in Wave II.
-To implement it, each unvalidated patient’s residual was calculated based on the naive healthcare utilization model estimates (Section 3.3) following (4), as
+Each unvalidated patient’s residual was calculated based on the naive healthcare utilization model estimates (Section 3.3) following (4) as
 $r_{i}^{*}=Y_{i}-\frac{1}{1+\exp\left\{-\left(-1.383+0.945X_{i}^{*}+0.103Z_{i}\right)\right\}}.$
-Then, patients were ranked by $r_{i}^{*}$, and the $24$ patients with the smallest residuals and the $24$ patients with the largest residuals were selected for validation.
+Then, patients were ranked by $r_{i}^{*}$.
+The $24$ patients with the smallest residuals and the $24$ patients with the largest residuals were selected for validation.
 Here, the RS design prioritized (i) patients with zero hospitalizations or ED visits but high predicted probabilities of having at least one and (ii) patients with at least one hospitalization or ED visit but high predicted probabilities of having zero.
-That is, patients whose observed $Y_{i}$ aligned well with the naive model’s prediction were not considered to be very informative and thus were deprioritized for validation.
+That is, patients whose $Y_{i}$ aligned with the naive model’s prediction were not considered to be very informative and were deprioritized for validation.
 2.6.2 Testing the Validation Protocol and Establishing Inter-Auditor Agreement
 The validation protocol evolved slightly throughout the waves, as well.
-After completing the first four patients’ chart reviews (Pilot), auditors and analysts met to discuss any protocol modifications that would be helpful before proceeding.
-For example, the auditor finding “extracted value not found” was not one of the original options, and the team discussed how these data points should be recorded.
-Adding an optional “Notes” column, where auditors could include any additional information, also resulted from this discussion.
-The Pilot also presented a valuable opportunity to evaluate inter-auditor agreement.
+After completing the Pilot, auditors and analysts met to discuss any protocol modifications that would be helpful.
+For example, the auditor finding “extracted value not found” was not originally an option, and we discussed how these data points should be recorded.
+Adding an optional auditor “Notes” column resulted from this conversation, too.
+We also evaluated inter-auditor agreement in the Pilot.
 We started with two auditors on our team.
-Each reviewed the four patients chosen for the Pilot, allowing us to calculate inter-auditor agreement in this doubly-validated sample.
-Later, two new auditors joined the team, each reviewing these four patients’ records as well.
+Each off them reviewed the four patients chosen for the Pilot, allowing us to calculate inter-auditor agreement in this doubly-validated sample.
+Later, two new auditors joined and also reviewed these four patients’ records.
 Ultimately, agreement in the four auditors’ findings was calculated using Fleiss’ Kappa [51], as detailed in the Supplemental Materials.
-We found high agreement between the auditors’ findings for labs ($\kappa=0.90$) and vitals ($\kappa=0.81$).
+We found high agreement for labs ($\kappa=0.90$) and vitals ($\kappa=0.81$).
 3 Results
 3.1 Cohort Description and Study Objectives
-We considered adult patients ($18$-$65$ years old) with at least one primary care outpatient encounter at Atrium Health Wake Forest Baptist Hospital.
+We considered adult patients (18186565 years old) with at least one primary care outpatient encounter at Atrium Health Wake Forest Baptist Hospital.
 The $1000$ sampled patients had a median age of $48$ years and a median income of \$52,320.
@@
 During the $2$-year study period, $318$ patients ($32\%$) had at least one hospitalization or ED visit.
-In unvalidated EHR data, error-prone ALI was right-skewed (Supplemental Figure S3).
-With a median of $0.33$ (interquartile range $=0.17,0.50$), patients experienced one out of every three non-missing components on their whole-person health, on average.
+In the unvalidated EHR data, error-prone ALI was right-skewed (Supplemental Figure S3).
+With a median of $0.33$ (interquartile range =0.17,0.50=0.17,0.50), patients experienced one out of every three non-missing components, on average.
 We validated ALI data for $100$ patients.
-First, we discuss the EHR data quality (Section 3.2).
-Second, we present the healthcare utilization model (Section 3.3).
-Within these sections, results are discussed chronologically, beginning with findings from the unvalidated EHR data and ending with those following the completion of all validation waves.
-We also discuss how these findings inform future implementations of the ALI in this healthcare system (Section 3.4).
+First, we discuss the EHR data quality.
+Second, we present the healthcare utilization model.
+Within these sections, results are introduced chronologically, beginning with findings from the unvalidated EHR data and ending with those based on all validation waves.
+We also explore how these findings inform future implementations of the ALI in this healthcare system.
 3.2 Validation Study Findings: EHR Data Quality
-As expected, the extracted EHR data had lots of missingness.
-Rare labs for inflammation components were missing for $>95\%$ of patients (Supplemental Figure S4A).
-Labs for metabolic components were missing for $21\%$ of patients.
-Common patterns of missing data are displayed in Supplemental Figure S5.
-Measurements for the cardiovascular components were non-missing for basically everyone.
+As expected, the extracted EHR data had lots of missingness (Supplemental Figure S4).
+Rare inflammation labs were missing for $95\%$ of patients (Supplemental Figure S5A).
+Metabolic labs were missing for $21\%$ of patients.
+Cardiovascular components (from vitals) were non-missing for basically everyone.
 Fortunately, most patients ($78\%$) had at least five non-missing ALI components (Supplemental Figure S5B).
-3.2.1 Pilot + Wave I Chart Reviews
+3.2.1 Pilot and Wave I Chart Reviews
 […]
-Figure 5: Auditors’ findings from the validation of A) $52$ patients’ data in the Pilot + Wave I, B) $48$ patients’ data in Wave II, and C) $100$ patients’ data in all waves combined.
+Figure 5: Auditors’ findings from the validation of A) $52$ patients’ data in the Pilot and Wave I, B) $48$ patients’ data in Wave II, and C) $100$ patients’ data in all waves combined.
 These findings refer to the original numeric measurements (before discretizing them into allostatic load index components), and there could be multiple per patient per variable.
-Auditor findings were mutually exclusive and specific to whether the measurements were originally non-missing (“extracted value correct,” “extracted value incorrect,” or “extracted value not found”) or missing (“no auxiliary information found” or “auxiliary information found”) in the extracted EHR data.
-Figure 6: Comparison of the discretized allostatic load index components from the unvalidated EHR data to A) $52$ patients’ data in the Pilot + Wave I, B) $48$ patients’ data in Wave II, and C) $100$ patients’ data in all waves combined.
+Auditor findings were mutually exclusive and specific to whether the measurements were originally non-missing (“extracted value correct,” “extracted value incorrect,” or “extracted value not found”) or missing (“no auxiliary information found” or “auxiliary information found”) in the extracted electronic health records data.
+Figure 6: Comparison of the discretized allostatic load index components from the unvalidated electronic health records (EHR) data to A) $52$ patients’ data in the Pilot and Wave I, B) $48$ patients’ data in Wave II, and C) $100$ patients’ data in all waves combined.
 Fewer vitals than labs matched Epic ($59\%$ versus $81\%$), with $31\%$ not found.
 We discovered that the EHR data extraction algorithm expected every patient encounter to have vitals.
-For encounters where vitals were not taken (e.g., physical therapy), values were imputed from other encounters; imputed values could not be found in Epic by the auditors.
-Thus, we uncovered an area for potential improvement in the current EHR data extraction protocol.
-Incorporating the audit roadmap into the validation protocol improved the completeness of the data.
+When vitals were not taken (e.g., physical therapy), values were imputed from other encounters; auditors could not find these values in Epic.
+Thus, we uncovered an improvement area in the EHR data extraction protocol.
+Incorporating the roadmap into the validation protocol improved the completeness of the data.
 The number of missing values decreased for nearly all ALI components (Supplemental Figure S5A).
 Homocysteine saw the biggest improvement, with auxiliary information located for $23$ of $52$ validated patients ($44\%$).
-The number of non-missing ALI components per patient increased from $6.6$ to $7.5$, on average.
-Overall, $48$ of $177$ missing components ($27\%$) were recovered through validation (Figure 6A).
+The median number of non-missing ALI components per patient increased from $7$ to $8$.
+Overall, $48$ of $177$ missing components ($27\%$) were recovered (Figure 6A).
 3.2.2 Wave II Chart Reviews
-In Wave II, $3867$ data points were validated ($718$ labs, $3149$ vitals), which was fewer than in the Pilot + Wave I.
-This difference was explained by fewer patient encounters, with $16$-$453$ validated data points per patient (median $=65.5$).
-The estimated error rates for non-missing components (TPR $=99\%$, FPR $=3\%$) and missing data recovery rate ($17\%$) in Wave II closely resembled the Pilot + Wave I (Figure 6B).
-Still, some auditors’ findings differed (Figure 5B).
-For example, more vitals matched Epic ($65\%$).
-Combined findings from all validation waves indicated that non-missing EHR data were of excellent quality (TPR $=99.5\%$, FPR $=2\%$), and auxiliary information could be located for $21\%$ of missing components (Figures 5C and 6C).
-On average, the number of non-missing components per patient increased from $6$ to $7$.
+In Wave II, $3867$ data points were validated ($718$ labs, $3149$ vitals), which was fewer than in the Pilot and Wave I.
+This difference was explained by fewer patient encounters, with 1616453453 validated data points per patient (median =65.5=65.5).
+The estimated error rates for non-missing components (TPR $99\%$, FPR $3\%$) and missing data recovery rate ($17\%$) in Wave II closely resembled the Pilot and Wave I (Figure 6B).
+Still, some auditors’ findings differed (Figure 5B), like more vitals matching Epic ($65\%$).
+Combined findings from all validation waves indicated that non-missing EHR data were of excellent quality (TPR $99.5\%$, FPR $2\%$), and auxiliary information was located for $21\%$ of missing components (Figures 5C and 6C).
+The median number of non-missing components per patient increased from $6$ to $7$.
 Supplemental Tables S1–S2 contain detailed findings.
 3.3 Analysis Findings: Whole-Person Health and Healthcare Utilization
-Using ALI values from the EHR data, we fit first the naive healthcare utilization model.
-The expected odds of engaging in the healthcare system for an $18$-year-old patient with an ALI of 00 were $0.25$ (95% confidence interval [95% CI]: $0.18$, $0.36$).
-For every $0.1$-point (one-component) increase in ALI, these odds were expected to increase by $10\%$ (OR $=1.10$, 95% CI: $1.03$, $1.18$), adjusting for age.
-For every $10$ years older a patient was, their odds were expected to increase by $11\%$ (OR $=1.11$, 95% CI: $1.00,1.23$), adjusting for ALI.
+Using the unvalidated EHR data, we first fit the naive model.
+The expected odds of engaging in the healthcare system for an $18$-year-old patient with an ALI of 0 were $0.25$ (95% confidence interval [95% CI]: $0.18$, $0.36$).
+For every $0.1$-point (one-component) increase in ALI, these odds were expected to increase by $10\%$ (OR =1.10=1.10, 95% CI: $1.03$, $1.18$), adjusting for age.
+For every $10$ years older a patient was, their odds were expected to increase by $11\%$ (OR =1.11=1.11, 95% CI: $1.00,1.23$), adjusting for ALI.
 While these findings align with the literature, we wanted to ensure that missingness and potential errors in the EHR data did not lead to spurious conclusions.
-3.3.1 Preliminary Estimates Based on the Pilot + Wave I Validation
-Incorporating the Pilot + Wave I, we re-fit the healthcare utilization model with the first $52$ patients’ validated data and all $1000$ patients’ unvalidated data.
-These “preliminary SMLEs” were used to design the simulations in Section 2.6.1.
-Here, we compare them to the naive model.
-For an 18-year-old patient with an ALI of 00, the expected odds of engaging in the healthcare system decreased to $0.17$ (95% CI: $0.12$, $0.26$).
-The association between whole-person health and healthcare utilization appeared even stronger, with an estimated 21% increase in odds for each $0.1$-point increase in ALI (OR $=1.21$, 95% CI: $1.17,1.25$), adjusting for age.
-The adjusted OR $=1.10$ for age was relatively unchanged (95% CI: $1.01,1.21$).
-Thus, the preliminary SMLEs reaffirmed that worse whole-person health was associated with higher expected odds of engaging in the healthcare system, perhaps more strongly than the naive model indicated.
+3.3.1 Preliminary Estimates Based on the Pilot and Wave I Validation
+Incorporating the Pilot and Wave I, we re-fit the model with the first $52$ patients’ validated data and all $1000$ patients’ unvalidated data.
+Simulations in Section 2.6.1 were designed from these “preliminary SMLEs.” For an 18-year-old patient with an ALI of 0, the expected odds of engaging in the healthcare system decreased to $0.17$ (95% CI: $0.12$, $0.26$).
+The association between ALI and healthcare utilization appeared even stronger, with an estimated 21% increase in odds for each $0.1$-point increase (OR =1.21=1.21, 95% CI: $1.17,1.25$), adjusting for age.
+The adjusted OR =1.10=1.10 for age was relatively unchanged (OR =1.10=1.10, 95% CI: $1.01,1.21$).
+Preliminary SMLEs reaffirmed that worse whole-person health was associated with higher expected odds of engaging in the healthcare system, perhaps more strongly than the naive model indicated.
 3.3.2 Final Estimates Based on All Waves of Validation
-After completing all three waves, we used the $100$ patients’ validated data with all $1000$ patients’ unvalidated data to re-fit the model one last time.
-All final estimates fell between the naive model and preliminary SMLEs.
-The expected odds of engaging in the healthcare system for an 18-year-old patient with an ALI of 00 were $0.23$ (95% CI: $0.16$, $0.33$).
-For each $0.1$-point increase in ALI, a $12\%$ increase in odds was expected (OR $=1.12$, 95% CI: $1.09,1.15$), adjusting for age.
-Again, older patients were expected to be more likely to engage, controlling for ALI (OR $=1.11$, 95% CI: $1.02,1.20$).
-Additional technical details, including R code, are available in the Supplemental Materials.
+After completing all three waves, we used the $100$ patients’ validated data with all $1000$ patients’ unvalidated data to re-fit the SMLEs one last time.
+These final estimates all fell between the naive model and preliminary SMLEs.
+The expected odds of engaging in the healthcare system for an 18-year-old patient with an ALI of 0 were $0.23$ (95% CI: $0.16$, $0.33$).
+For each $0.1$-point increase in ALI, a $12\%$ increase in odds was expected (OR =1.12=1.12, 95% CI: $1.09,1.15$), adjusting for age.
+Again, older patients were expected to be more likely to engage (adjusted OR =1.11=1.11, 95% CI: $1.02,1.20$).
+Technical details and R code are in the Supplemental Materials.
 3.4 Informing Implementation in the Learning Health System
-Ultimately, all models agreed that worse whole-person health (i.e., higher ALI) was associated with a higher likelihood of engaging in the healthcare system, adjusting for age.
-Chart reviews uncovered very high quality of non-missing EHR data and located auxiliary information for some missing data.
-Without validation, knowing whether we could trust the EHR data and whether the EHR-derived ALI was still a valid computable phenotype for whole-person health would not have been possible.
-Thus, while the models led to the same conclusions here, validating EHR data before using them in practice or research is critical.
+All models agreed that worse whole-person health (higher ALI) was associated with a higher likelihood of engaging in the healthcare system, adjusting for age.
+Chart reviews uncovered high quality of non-missing EHR data and located auxiliary information for some missing data.
+Without validation, knowing whether we could trust the EHR data and whether the EHR-derived ALI was a valid computable phenotype for whole-person health would not have been possible.
+Thus, while the models led to the same conclusions, validating EHR data before using them in practice or research is critical.
 Our findings can now inform implementation strategies to incorporate the ALI into this learning health system.
-Having established ALI as a strong predictor of healthcare utilization, embedding it in the EHR is a critical next step.
-If ALI automatically prepopulated in patients’ electronic charts, clinicians would have easy access to it when making care decisions and researchers could extract it for further studies.
-This healthcare system previously implemented another score, the EFI, across the patient population aged $65+$.
-Building on that experience and encouraged by the ALI’s predictive ability, we plan to operationalize it for patients $18$ years and older.
+For example, if the ALI automatically prepopulates in patients’ charts, clinicians would have easy access to it when making care decisions, and researchers could extract it for further studies.
+This healthcare system previously implemented another score, the EFI, across the patient population aged $65$ and older.
+Building on that experience and encouraged by the ALI’s predictive ability, we plan to operationalize it for patients aged $18$ and older.
 Of course, we want to embed the validated ALI, and only a small subset of patients can undergo chart reviews.
-Fortunately, in obtaining SMLEs for the healthcare utilization model, we robustly estimated the exposure error mechanism, $\Pr(X|X^{*},Z)$.
+Fortunately, in obtaining SMLEs for the healthcare utilization model, we robustly estimate the exposure error mechanism, $\Pr(X|X^{*},Z)$.
 When scaling up across the healthcare system, the estimated exposure error mechanism can be used to efficiently predict patients’ validated ALI from unvalidated EHR data (Figure 7).
 See the Supplemental Materials for details on calculating these predictions.
-Figure 7: Among patients with chart reviews (triangles), validated allostatic load index (ALI) could differ noticeably from the unvalidated version using the EHR data.
+Figure 7: Among patients with chart reviews (triangles), validated allostatic load index (ALI) could differ noticeably from the unvalidated version using the unvalidated electronic health records (EHR) data.
 For patients without chart reviews (squares), validated ALI could be robustly predicted using the estimated exposure error mechanism.
@@
 Predicting healthcare utilization is a priority in the learning health system, with important implications at both patient and institutional levels.
-Establishing standardized, valid measures of whole-person health (like ALI) that are associated with utilization and can scale up to the entire institution is an integral step.
+Establishing standardized, valid measures of whole-person health (like the ALI) that are associated with utilization and can scale up to the entire institution is an integral step.
 While EHR data contain variables to calculate these measures, missing data is a key challenge, and non-missing data are still prone to errors.
@@
 Also, existing validation protocols do not overcome missingness.
-We propose novel methods to maximize information gained from a partial validation study of $100$ patients from an EHR dataset of $1000$ through the statistical model, validation protocol, and targeted strategy to select patients.
+We propose novel methods to maximize information gained from a partial validation study through the statistical model, validation protocol, and targeted patient selection strategy.
 We focused on estimating the association between ALI and healthcare utilization.
-Partial validation can also efficiently improve estimates of the ALI itself, in preparation for implementation in the EHR.
+In preparation for EHR implementation, partial validation can also efficiently improve estimates of the ALI itself.
 The proposed analytical and design strategies were devised to ensure robust and efficient estimates.
-Specifically, the statistical model avoids making restrictive assumptions about the error mechanism, relating the unvalidated and validated ALI, while incorporating all available information to achieve precise estimates (i.e., low variance).
-Moreover, the targeted design strategies are tailored to the preliminary validation data and chosen statistical model, selecting patients most informative for the clinical question.
+Specifically, the statistical model avoids making restrictive assumptions about the error mechanism, while incorporating all available information to achieve precise estimates (i.e., low variance).
+Moreover, the targeted design strategies are tailored to the preliminary validation data and chosen statistical model, selecting patients who are most informative for the clinical question.
 Our enriched validation protocol builds on the literature (e.g., [52]) by incorporating a clinically-driven “roadmap” to locate auxiliary information about missing values.
-The robust SMLEs can handle conditioning on error-prone $X^{*}$ and maybe one other covariate $Z$ in the error mechanism, due to the curse of dimensionality.
+Due to the curse of dimensionality, the SMLEs can handle conditioning on error-prone $X^{*}$ and maybe one other covariate $Z$ in the error mechanism.
 We assumed that validated ALI depended only on its unvalidated version.
-However, if a more complicated error structure is expected, data pre-processing will be needed before using the SMLEs Tao2021.
-While our roadmap helped recover some missing ALI components, validated patients could still have incomplete data.
+However, if a more complicated error structure is expected, data pre-processing will be needed before using the SMLEs [53].
+While our roadmap recovered some missing ALI components, validated patients could still have incomplete data.
 Clinical guidelines were only defined for missing values presumed to be unhealthy.
-If auditors were given additional guidance on what auxiliary information might mean that a missing value is healthy, missingness could be further reduced.
-Others have described missingness in the EHR as selection bias and sought to correct for it using approaches like IPW [53].
+Missingness could be further reduced if auditors were given additional guidance on auxiliary information for healthy values.
+Others have described missingness in the EHR as selection bias and corrected for it using approaches like IPW [54].
 Applying such approaches to our post-validation data may reduce the impact of residual missingness.
-Alternatively, we could adopt other allostatic load calculations, like the index of cardiometabolic health (ICMH), which rely less on rare labs and may face fewer missing components [9].
+Alternatively, we could adopt other allostatic load calculations, like the index of cardiometabolic health, which rely less on rare labs and may face fewer missing components [8].
 EHR data are susceptible to “informative presence bias,” wherein the process through which patients engage in the healthcare system can bias associations [55, 56, 57].
 Our decision to adapt the Seeman et al.
-[7] definition of the ALI as the proportion of non-missing components could be considered a summary measure.
+(2001) [9] ALI as the proportion of non-missing components could be considered a summary measure.
 A recent review found summary measures to be the most common approach to handling informative presence bias [58].
-Whole-person health scores like ALI are among many computable phenotypes currently applied to EHR data [16].
-Other phenotyping algorithms, including EHRshot [58], also seek to translate observable characteristics into measures of underlying patient states.
-Validation is a critical step in their development and typically involves chart reviews to obtain some “gold standard” labels [59, 60].
+The ALI is among many computable phenotypes currently applied to EHR data [16].
+Other phenotyping algorithms, including EHRshot [59], translate observable characteristics into measures of underlying patient states.
+Validation is a critical step in their development and typically involves chart reviews to obtain “gold standard” labels [60, 61].
 When phenotyping with machine learning methods, this validation set generally needs to be an SRS.
-Meanwhile, our statistical modeling approach accommodates targeted sampling of the validation study.
-This flexibility creates opportunities to promote the efficiency of downstream analyses (e.g., associations between the phenotype and suspected risk factors).
+Meanwhile, our statistical model accommodates targeted validation sampling, creating opportunities to promote the efficiency of downstream analyses (e.g., associations between the phenotype and suspected risk factors).
 There are several interesting directions for future work.
 First, a longitudinal approach to modeling healthcare utilization, wherein patients’ ALI can vary over time, might provide more information.
-Second, auditors recorded their findings in a Microsoft Excel workbook.
-For ease of use and accuracy, an electronic REDCap [61] interface could be built following published guidelines [35].
-Third, targeted validation strategies could be used to improve predictive ability, rather than efficiency [62, 63].
+Second, auditors recorded their findings in Microsoft Excel, but an electronic REDCap [62] interface would improve usability and accuracy [34].
+Third, targeted validation strategies could be used to improve predictive ability rather than efficiency [63, 64].
 5 Conclusion
@@
 However, errors and missingness in these data pose challenges in verifying and operationalizing such measures.
-Validation of even a subset of patients’ data is time- and cost-intensive, but provides crucial insights into data quality.
+Validation of even a subset of patients’ data is time- and cost-intensive but provides crucial insights into data quality.
 Statistical and biomedical informatics methods can maximize information gained from such partial validation.